### v2

# On properties and numerical computation of critical points of eigencurves of bivariate matrix pencils

Abstract
We investigate critical points of eigencurves of bivariate matrix pencils $A+\lambda B+\mu C$. Points $(\lambda,\mu)$ for which $\det(A+\lambda B+\mu C)=0$ form algebraic curves in $\mathbb{C}^{2}$ and we focus on points where $\mu^{\prime}(\lambda)=0$. Such points are referred to as zero-group-velocity (ZGV) points, following terminology from engineering applications. We provide a general theory for the ZGV points and show that they form a subset (with equality in the generic case) of the 2D points $(\lambda_{0},\mu_{0})$, where $\lambda_{0}$ is a multiple eigenvalue of the pencil $(A+\mu_{0}C)+\lambda B$, or, equivalently, there exist nonzero $x$ and $y$ such that $(A+\lambda_{0}B+\mu_{0}C)x=0$, $y^{H}(A+\lambda_{0}B+\mu_{0}C)=0$, and $y^{H}Bx=0$.
We introduce three numerical methods for computing 2D and ZGV points. The first method calculates all 2D (ZGV) points from the eigenvalues of a related singular two-parameter eigenvalue problem. The second method employs a projected regular two-parameter eigenvalue problem to compute either all eigenvalues or only a subset of eigenvalues close to a given target. The third approach is a locally convergent Gauss–Newton-type method that computes a single 2D point from an inital approximation, the later can be provided for all 2D points via the method of fixed relative distance by Jarlebring, Kvaal, and Michiels. In our numerical examples we use these methods to compute 2D-eigenvalues, solve double eigenvalue problems, determine ZGV points of a parameter-dependent quadratic eigenvalue problem, evaluate the distance to instability of a stable matrix, and find critical points of eigencurves of a two-parameter Sturm-Liouville problem.
keywords:
zero-group-velocity point, 2D point, bivariate matrix pencil, singular two-parameter eigenvalue problem, 2D-eigenvalue, double eigenvalue, distance to instability, two-parameter Sturm-Liouville problem
pacs:
pacs:
MSC Classification]65F15, 65F50, 15A18, 15A22
Funding]The work was supported by the Slovenian Research and Innovation Agency (grants N1-0154 and P1-0294).
1 Introduction
We consider parameter-dependent linear eigenvalue problems of the form
$(A+\lambda B+\mu C)x=0,$ (1)
where $A,B,C\in\mathbb{C}^{n\times n}$, $\lambda,\mu\in\mathbb{C}$ and $x\in\mathbb{C}^{n}$ is nonzero. We are interested in properties and numerical computation of points $(\lambda_{0},\mu_{0})\in\mathbb{C}^{2}$ such that $\mu$ is an analytic function of $\lambda$, $\mu_{0}=\mu(\lambda_{0})$, and $\mu^{\prime}(\lambda_{0})=0$. We will show that such points belong to a larger set of points where $\lambda_{0}$ is a multiple eigenvalue of the pencil $(A+\mu_{0}C)+\lambda B$. Such points are related to problems in applied mathematics and engineering, e.g., in elastodynamics [19], the double eigenvalue problem [18, 25], the distance to instability of a stable matrix [8], and the 2D eigenvalue problem [32].
To ensure that the pencil (1) behaves regularly with respect to both parameters we assume that the problem is biregular according to the following definition.
Definition 1.1.
A bivariate matrix pencil $A+\lambda B+\mu C$ is biregular if for each $(\lambda_{0},\mu_{0})\in\mathbb{C}^{2}$ the generalized eigenvalue problems (GEPs)
$\big((A+\lambda_{0}B)+\mu C\big)x=0$ (2)
and
$\big((A+\mu_{0}C)+\lambda B\big)x=0$ (3)
are both regular, i.e., $\det(A+\lambda_{0}B+\mu C)\not\equiv 0$ and $\det(A+\lambda B+\mu_{0}C)\not\equiv 0$.
It is easy to see that a bivariate pencil $A+\lambda B+\mu C$ is biregular if and only if the characteristic polynomial $p(\lambda,\mu):=\det(A+\lambda B+\mu C)$ does not have a divisor of the form $\lambda-\lambda_{0}$ or $\mu-\mu_{0}$. A sufficient condition for biregularity is that matrices $B$ and $C$ are both nonsingular. If $A+\lambda B+\mu C$ is biregular then for each $\lambda_{0}\in\mathbb{C}$ there are $n$ eigenvalues $\mu\in\mathbb{C}\cup\{\infty\}$ of (2). As a result, the set of points $(\lambda,\mu)\in\mathbb{C}^{2}$ such that $\det(A+\lambda B+\mu C)=0$ is composed of algebraic curves in $\mathbb{C}^{2}$ that are called eigencurves. If $(\lambda_{0},\mu_{0})\in\mathbb{C}^{2}$ is a point on an eigencurve such that $\mu_{0}$ is a simple eigenvalue of the GEP (2), then we can locally parameterize $\mu$ as an analytic function of $\lambda$ such that $\mu(\lambda_{0})=\mu_{0}$ and define $\mu^{\prime}(\lambda_{0})$.
Theorem 1.2.
Let $\mu_{0}\in\mathbb{C}$ be a simple eigenvalue of a regular GEP $((A+\lambda_{0}B)+\mu C)x=0$. Then there exist analytic functions $\mu(\lambda)$ and $x(\lambda)\neq 0$ in a neighbourhood of $\lambda_{0}$ such that $\big(A+\lambda B+\mu(\lambda)C\big)x(\lambda)=0$ and $\mu(\lambda_{0})=\mu_{0}$.
Proof.
We know that the theorem holds for the standard eigenvalue problem where $C=I$, see, e.g., [11, Thm. 2]. Since $(A+\lambda_{0}B)+\mu C$ is a regular GEP, it follows that the matrix $M(\alpha):=\alpha(A+\lambda_{0}B)+C$ is nonsingular for almost all $\alpha\in\mathbb{C}$. It is easy to see that $\big((A+\lambda_{0}B)+\mu C\big)x=0$ if and only if $((A+\lambda_{0}B)+\eta M(\alpha))x=0$ for $\eta=\mu/(1-\alpha\mu)$. As the theorem holds for the standard eigenvalue problem $(M(\alpha)^{-1}(A+\lambda_{0}B)+\eta I)x=0$, it thus holds for the GEP (2) as well. ∎
Definition 1.3.
Let $(\lambda_{0},\mu_{0})\in\mathbb{C}^{2}$ be such that $\mu_{0}$ is a simple eigenvalue of a regular GEP $(A+\lambda_{0}B)+\mu C)x=0$ and let $\mu(\lambda)$ be an analytic function in a neighbourhood of $\lambda_{0}$ such that $\det(A+\lambda B+\mu(\lambda)C)=0$ and $\mu(\lambda_{0})=\mu_{0}$. We say that $(\lambda_{0},\mu_{0})$ is a ZGV (zero-ground-velocity) point of the bivariate pencil $A+\lambda B+\mu C$ if $\mu^{\prime}(\lambda_{0})=0$.
The expression ZGV point comes from the study of waves in elastodynamics, where the angular frequency $\omega$ of a guided wave is related to the wavenumber $k$ via a dispersion relation $\omega(k)$. Of special interest are zero-group-velocity (ZGV) points $(k_{*},\omega_{*})$ on the dispersion curves, where the group velocity $c_{g}=\frac{\partial\omega}{\partial k}$ of a wave vanishes whereas the wave number $k_{*}$ remains finite, see, e.g., [29].
The main objective of this work is to clarify the algebraic structure underlying the ZGV points and design numerical methods that can compute all such points. The key for this will be the characterization of the ZGV points as eigenvalues of an associated singular two-parameter eigenvalue problem.
The rest of this work is structured as follows. In Section 2 we discuss the properties of ZGV points, introduce 2D points, and show that each ZGV point $(\lambda_{0},\mu_{0})$ is also a 2D point with $\lambda_{0}$ being a multiple eigenvalue of the GEP (3). In Section 3 we review two-parameter eigenvalue problems (2EPs) and singular GEPs, both will be essential tools in later sections. Section 4 contains our main result that 2D points are eigenvalues of a singular 2EP, which we exploit to derive a numerical method for computing all 2D and ZGV points. In Subsection 4.1 we show that the related singular 2EP can be solved more efficiently by projecting it into a nonsingular 2EP and applying the numerical method from [15]. In Section 5 we provide a Gauss–Newton-type method that computes a 2D point from a good initial approximation and prove its quadratic convergence in the generic case. Approximate solutions for 2D points can be obtained by the method of fixed relative distance [18] that is presented in Section 6. Finally, Section 7 explores several applications of 2D and ZGV points, illustrated with numerical examples.
2 Main properties
The next auxiliary lemma contains well-known results (see, e.g., [9, 31]) on eigenvalues and eigenvectors of regular GEPs that we will use frequently in the following.
Lemma 2.1.
Let $\lambda_{0}\in\mathbb{C}$ be a finite eigenvalue of a regular GEP $(A+\lambda B)x=0$.
  1. [left=0em]
  2. a)
    If $\lambda_{0}$ is simple and $x_{0}$ and $y_{0}$ are the corresponding right and left eigenvector, then $y_{0}^{H}Bx_{0}\neq 0$.
  3. b)
    $\lambda_{0}$ is simple if and only if it is geometrically simple and $y_{0}^{H}Bx_{0}\neq 0$, where $x_{0}$ and $y_{0}$ are the corresponding right and left eigenvector.
  4. c)
    If algebraic multiplicity of $\lambda_{0}$ is larger than the geometric multiplicity, then there exist nonzero vectors $x_{0}$ (an eigenvector) and $z_{0}$ (a generalized eigenvector of degree two), such that
    $(A+\lambda_{0}B)x_{0}=0\quad\textrm{and}\quad(A+\lambda_{0}B)z_{0}+Bx_{0}=0.$
The following lemma gives necessary conditions for a ZGV point.
Lemma 2.2.
If $(\lambda_{0},\mu_{0})\in\mathbb{C}^{2}$ is a ZGV point of a bivariate pencil $A+\lambda B+\mu C$ and $x_{0}$ and $y_{0}$ are the corresponding right and left eigenvector, then
$y_{0}^{H}Bx_{0}=0$ (4)
and
$y_{0}^{H}Cx_{0}\neq 0.$ (5)
Proof.
Let $\mu(\lambda)$ and $x(\lambda)$ be analytic functions from Theorem 1.2 such that $\big(A+\lambda B+\mu(\lambda)C\big)x(\lambda)=0$, $x(\lambda_{0})=x_{0}$, and $\mu(\lambda_{0})=\mu_{0}$. By differentiating we get
$\big(A+\lambda B+\mu(\lambda)C\big)x^{\prime}(\lambda)+\big(B+\mu^{\prime}(\lambda)C\big)x(\lambda)=0.$ (6)
Multiplying (6) from the left by $y_{0}^{H}$ and taking into account that at a ZGV point $u^{\prime}(\lambda_{0})=0$ yields the condition (4). The condition (5) follows from Lemma 2.1 from the demand that $\mu_{0}$ is a simple eigenvalue of the GEP (2). ∎
Based on the condition (4) from Lemma 2.2 we introduce a larger set of 2D points for the bivariate pencil $A+\lambda B+\mu C$. The name comes from a relation to the 2D-eigenvalue problem [21] that we will explore later in Subsection 7.1. We remark that for the symmetric matrices and a simple eigenvalue $\mu_{0}$ the necessary condition (4) for $(\lambda_{0},\mu_{0})$ to be a critical point on an eigencurve was noticed in [27].
Definition 2.3.
We say that $(\lambda_{0},\mu_{0})\in\mathbb{C}^{2}$ is a 2D point of a bivariate pencil $A+\lambda B+\mu C$ if there exist nonzero vectors $x_{0}$ and $y_{0}$ such that
$\begin{split}(A+\lambda_{0}B+\mu_{0}C)x_{0}&=0,\\ y_{0}^{H}(A+\lambda_{0}B+\mu_{0}C_{0})&=0,\\ y_{0}^{H}Bx_{0}&=0.\end{split}$ (7)
It follows from Lemma 2.2 that ZGV points are a subset of 2D points. In the following we will identify which 2D points are also ZGV points.
Lemma 2.4.
Let the GEP (3) be regular for $\mu_{0}\in\mathbb{C}$. Then $(\lambda_{0},\mu_{0})$ is a 2D point of the bivariate pencil $A+\lambda B+\mu C$ if and only if $\lambda_{0}\in\mathbb{C}$ is a multiple eigenvalue of (3).
Proof.
Suppose that $\lambda_{0}$ is a simple eigenvalue of (3) and $x_{0}$ and $y_{0}$ are the corresponding right and left eigenvector. Then $y_{0}^{H}Bx_{0}\neq 0$ by Lemma 2.1, which contradicts the condition from Definition 2.3.
For a proof in the other direction, let $\lambda_{0}$ be a multiple eigenvalue of (3). We consider two options.
  1. [left=0em]
  2. a)
    $\lambda_{0}$ is semisimple, i.e., the geometric multiplicity of $\lambda_{0}$ equals the algebraic multiplicity. In this case the left and the right eigenvector belong to subspaces of dimension at least two and there exist nonzero $x_{0}\in\mathop{\rm Ker}\nolimits(A+\mu_{0}C+\lambda_{0}B)$ and $y_{0}\in\mathop{\rm Ker}\nolimits((A+\mu_{0}C+\lambda_{0}B)^{H})$ such that $y_{0}^{H}Bx_{0}=0$.
  3. b)
    The algebraic multiplicity of $\lambda_{0}$ is larger than the geometric multiplicity. Then there exist nonzero vectors $x_{0}$ (an eigenvector) and $z$ (a generalized eigenvector of degree two) such that
    $\begin{split}(A+\lambda_{0}B+\mu_{0}C)x_{0}&=0,\\ (A+\lambda_{0}B+\mu_{0}C_{0})z+Bx_{0}&=0.\end{split}$ (8)
    By multiplying the second equation of (8) by the left eigenvector $y_{0}$, we get $y_{0}^{H}Bx_{0}=0$.
In both cases we get vectors $x_{0}$ and $y_{0}$ such that (7) holds and $(\lambda_{0},\mu_{0})$ is a 2D point. ∎
Remark 2.5.
Let $(\lambda_{0},\mu_{0})\in\mathbb{C}^{2}$ be a 2D point of a biregular bivariate pencil $A+\lambda B+\mu C$. If we denote by $a_{m}$ and $g_{m}$ the algebraic and geometric multiplicity of $\lambda_{0}$ as an eigenvalue of (3), respectively, then the following options are possible:
  1. [left=0em]
  2. a)
    $a_{m}\geq 2$, $g_{m}=1$, and $y_{0}^{H}Cx_{0}\neq 0$,
  3. b)
    $a_{m}\geq 2$, $g_{m}=1$, and $y_{0}^{H}Cx_{0}=0$,
  4. c)
    $a_{m}>g_{m}\geq 2$,
  5. d)
    $a_{m}=g_{m}\geq 2$,
where $x_{0}$ and $y_{0}$ are the right and the left eigenvector for the eigenvalue $\lambda_{0}$ of (3). The generic situation is a). In case d) $\lambda_{0}$ is a semisimple eigenvalue of (3), and a non-semisimple eigenvalue in cases a), b), c).
Throughout the paper we use the term generic as follows: a set $\mathcal{A}\subseteq\mathbb{C}^{m}$ is algebraic if it is the common zero set of finitely many complex polynomials in $m$ variables. A set $\mathcal{U}$ is called generic if its complement is contained in a proper algebraic set. Similarly, a property is generic in $\mathcal{U}\subseteq\mathbb{C}^{m}$, if there exists an algebraic set $\mathcal{A}\subseteq\mathbb{C}^{m}$ such that $\mathcal{U}\not\subseteq\mathcal{A}$ and the property holds on $\mathcal{U}\backslash(\mathcal{A}\cap\mathcal{U})$.
In line with Remark 2.5 we classify 2D points into points of type a), b), c), and d). We can now show that ZGV points of a biregular bilinear pencil are exactly 2D points of type a).
Theorem 2.6.
Let $(\lambda_{0},\mu_{0})\in\mathbb{C}^{2}$ be a 2D point of a biregular bivariate pencil $A+\lambda B+\mu C$. Then $(\lambda_{0},\mu_{0})$ is a ZGV point if and only if the geometric multiplicity of $\lambda_{0}$ as an eigenvalue of (3) is one and $y_{0}^{H}Cx_{0}\neq 0$, where $x_{0}$ and $y_{0}$ are the right and the left eigenvector.
Proof.
Let $(\lambda_{0},\mu_{0})$ be a ZGV point. As the geometric multiplicity of $\lambda_{0}$ as an eigenvalue of (3) is equal to the geometric multiplicity of $\mu_{0}$ as an eigenvalue of (2), cases c) and d) from Remark 2.5 are not possible. It follows from the condition (5) from Lemma 2.2 that $(\lambda_{0},\mu_{0})$ is a 2D point of type a).
For a proof in the other direction, let $(\lambda_{0},\mu_{0})$ be a 2D point of type a). Using the same argument as above we see that the geometric multiplicity of $\mu_{0}$ as an eigenvalue of (2) is one. The eigenvectors $x_{0}$ and $y_{0}$ are then uniquely defined up to a multiplication by a scalar and it follows from $y_{0}^{H}Cx_{0}\neq 0$ that $\mu_{0}$ is algebraically simple. Therefore, there exist analytic functions $\mu(\lambda)$ and $x(\lambda)$ from Theorem 1.2 such that $\big(A+\lambda B+\mu(\lambda)C\big)x(\lambda)=0$, $x(\lambda_{0})=x_{0}$, and $\mu(\lambda_{0})=\mu_{0}$. If we differentiate the equation as in the proof of Lemma 2.2 and multiply the derivative (6) from the left by $y_{0}^{H}$, we obtain
$y_{0}^{H}(B+\mu^{\prime}(\lambda_{0})C)x_{0}=0.$
Since $(\lambda_{0},\mu_{0})$ is a 2D point, $y_{0}^{H}Bx_{0}=0$, and, since $y_{0}^{H}Cx_{0}\neq 0$, it follows that $\mu^{\prime}(\lambda_{0})=0$ and $(\lambda_{0},\mu_{0})$ is a ZGV point. ∎
We will use the next result in numerical methods to detect if a computed 2D point is a ZGV point.
Corollary 2.7.
Let $(\lambda_{0},\mu_{0})\in\mathbb{C}^{2}$ be a 2D point of a biregular bivariate pencil $A+\lambda B+\mu C$. Then $(\lambda_{0},\mu_{0})$ is a ZGV point if and only if $\mu_{0}$ is a simple eigenvalue of the GEP (2).
Proof.
If $\mu_{0}$ is a simple eigenvalue of (2) and $x_{0}$ and $y_{0}$ are the right and the left eigenvector, then $y_{0}^{H}Cx_{0}\neq 0$. Another thing that follows from $\mu_{0}$ being simple is that the geometric multiplicity of $\lambda_{0}$ as an eigenvalue of (3) is one. Thus, $(\lambda_{0},\mu_{0})$ is a 2D point of type a) and it follows from Theorem 2.6 that $(\lambda_{0},\mu_{0})$ is a ZGV point.
A proof in the other direction follows directly from the definition of the ZGV point, where we require that $\mu_{0}$ is a simple eigenvalue of (2). ∎
The following two small examples illustrate the definition of ZGV and 2D points.
Example 2.8.
If we take
$A=\left[\begin{matrix}1&2&3&0\cr 2&0&1&0\cr 3&1&1&0\cr 0&0&0&-3\end{matrix}\right],\quad B=\left[\begin{matrix}1&0&1&0\cr 0&1&1&0\cr 1&1&0&0\cr 0&0&0&-3\end{matrix}\right],\quad C=\left[\begin{matrix}2&1&0&0\cr 1&3&0&0\cr 0&0&1&0\cr 0&0&0&1\end{matrix}\right],$ (9)
then the matrix $A+\lambda_{0}B$ is symmetric for $\lambda_{0}\in\mathbb{R}$ and, since $C$ is symmetric positive definite, all eigenvalues of $(A+\lambda_{0}B)+\mu C$ are real. The left hand side of Figure 1 shows the real eigencurves of $A+\lambda B+\mu C$ with the four real ZGV points (2.2645,1.3475)(-2.2645,-1.3475), (1.8172,0.17299)(-1.8172,-0.17299), $(0.28896,0.28248)$, and $(0.38688,1.7975)$. In addition, the problem has two complex ZGV points $(-10.4081\pm 3.8258i,7.7647\mp 2.9511i)$. There are additional three 2D points of type d) that are not ZGV points. These are the intersections of eigencurves at (1.5330,1.5991)(-1.5330,-1.5991), (1,0)(-1,0), and (0.3565,1.9305)(-0.3565,1.9305).
Figure 1: Real eigencurves with ZGV (red) and 2D (white) points of Examples 2.8 (left) and 2.9 (right).
Example 2.9.
If we take
$A+\lambda B+\mu C=\left[\begin{matrix}\lambda+\mu&1\cr 0&\lambda+2\mu\end{matrix}\right],$
then $(0,0)$ is the only 2D point. It is of type b) and the pencil does not have any ZGV points. For the corresponding eigencurves see the right hand side of Figure 1. Note that although the 2D point is the intersection of eigencurves like in the previous example, it is of a different type than the 2D points from Example 2.8.
It follows from Corollary 2.7 that ZGV points are exactly the subset of 2D points of $A+\lambda B+\mu C$ such that $\mu_{0}$ is a simple eigenvalue of $\big((A+\lambda_{0}B)+\mu C\big)x=0$. Based on Theorem 2.6 and Corollary 2.7 we can build criteria to identify 2D points that are also ZGV points, which means that all numerical methods for the computation of 2D points can be used to compute ZGV points as well.
The problem of finding 2D points is by Lemma 2.4 equivalent to the problem of finding values $\mu_{0}$ such that the GEP (3) has a multiple eigenvalue $\lambda_{0}$. Generically, see, e.g., [25], if the GEP (3) has a multiple eigenvalue $\lambda_{0}$, then this eigenvalue is a derogatory double eigenvalue, i.e., the algebraic multiplicity of $\lambda_{0}$ is two and the geometric multiplicity is one, thus it is a 2D point of type a) or b). In such case, a 2D point is a ZGV point if and only if it satisfies the condition (5). As generically this condition is always satisfied, the 2D points and ZGV points agree in the generic case, see also Theorem A.1.
We know that an $n\times n$ generic bivariate matrix pencil $A+\lambda B+\mu C$ has n(n1)n(n-1) 2D points [25], as this is the number of values $\mu$ such that a matrix $B^{-1}A+\mu B^{-1}C$ has a double eigenvalue (note that a generic $B$ is nonsingular). It follows that a biregular bivariate matrix pencil has n(n1)n(n-1) or less ZGV points, with an equality in the generic case when 2D points and ZGV points agree.
The numerical methods that we will present later are based on the theory on two-parameter eigenvalue problems and singular generalized eigenvalue problems that we introduce in the next section.
3 Auxiliary results
A two-parameter eigenvalue problem (2EP) has the form
$\begin{split}W_{1}(\lambda,\mu)x_{1}&:=(A_{1}+\lambda B_{1}+\mu C_{1})x_{1}=0,\\ W_{2}(\lambda,\mu)x_{2}&:=(A_{2}+\lambda B_{2}+\mu C_{2})x_{2}=0,\end{split}$ (10)
where $A_{i},B_{i},C_{i}\in\mathbb{C}^{n_{i}\times n_{i}}$, $x_{1},x_{2}$ are nonzero vectors and $\lambda,\mu\in\mathbb{C}$. If $(\lambda,\mu)$ and nonzero $x_{1},x_{2}$ satisfy (10), then $(\lambda,\mu)$ is an eigenvalue and $x_{1}\otimes x_{2}$ is the corresponding (right) eigenvector. Similarly, $y_{1}\otimes y_{2}$ is the corresponding left eigenvector if $y_{i}\neq 0$ and $y_{i}^{H}W_{i}(\lambda,\mu)=0$ for $i=1,2$.
Note that (1) has a form of one equation from (10). In the generic case its solutions $(\lambda,\mu)$, where $\det(A+\lambda B+\mu C)=0$, lie on curves in $\mathbb{C}^{2}$ and there are infinitely many such points. On the other hand, eigenvalues of a 2EP (10) are the intersections of curves
$\begin{split}p_{1}(\lambda,\mu)&:=\det(W_{1}(\lambda,\mu))=0,\\ p_{2}(\lambda,\mu)&:=\det(W_{2}(\lambda,\mu))=0.\end{split}$ (11)
In the generic case the curves $p_{1}(\lambda,\mu)=0$ and $p_{2}(\lambda,\mu)=0$ do not have a nontrivial common factor and by Bézout’s theorem (see, e.g., [30]) it follows that (10) has $n_{1}n_{2}$ eigenvalues.
A 2EP (10) is related to a coupled pair of GEPs
$\begin{split}\Delta_{1}z&=\lambda\Delta_{0}z,\\ \Delta_{2}z&=\mu\Delta_{0}z,\end{split}$ (12)
where $n_{1}n_{2}\times n_{1}n_{2}$ matrices
$\begin{split}\Delta_{0}&=B_{1}\otimes C_{2}-C_{1}\otimes B_{2},\\ \Delta_{1}&=C_{1}\otimes A_{2}-A_{1}\otimes C_{2},\\ \Delta_{2}&=A_{1}\otimes B_{2}-B_{1}\otimes A_{2}\end{split}$ (13)
are called operator determinants and $z=x_{1}\otimes x_{2}$ is a decomposable vector, for details see, e.g., [2]. A generic 2EP (10) is nonsingular, i.e., the corresponding operator determinant $\Delta_{0}$ is nonsingular. In this case (see, e.g., [2]), the matrices $\Delta_{0}^{-1}\Delta_{1}$ and $\Delta_{0}^{-1}\Delta_{2}$ commute and the eigenvalues of (10) agree with the joint eigenvalues of (12). By exploting this relation it is possible to numerically solve a nonsingular 2EP using standard tools for GEPs, see, e.g., [13].
If all linear combinations of $\Delta_{0}$, $\Delta_{1}$ and $\Delta_{2}$ are singular, we have a singular 2EP that is much more difficult to solve. It can still have a finite number of eigenvalues that can be computed by a staircase type algorithm from [24] or by an algorithm for singular GEPs from [14, 15]. In the singular case, the relation between problems (10) and (12) is much less understood, for details see [23, 20].
We say that a matrix pencil $A-\lambda B$, where $A,B\in\mathbb{C}^{n\times n}$, is singular if $\det(A-\lambda B)\equiv 0$. This is equivalent to the fact that the normal rank of the pencil, defined as
$\textrm{nrank}(A-\lambda B):=\max_{\xi\in\mathbb{C}}(\mathop{\rm rank}\nolimits(A-\xi B)),$
is strictly smaller than $n$. In such case $\lambda_{0}\in\mathbb{C}$ is a finite eigenvalue of $A-\lambda B$ if $\mathop{\rm rank}\nolimits(A-\lambda_{0}B)<\textrm{nrank}(A-\lambda B)$ and $\lambda_{0}=\infty$ is an eigenvalue if $\mathop{\rm rank}\nolimits(B)<\textrm{nrank}(A-\lambda B)$. Standard numerical methods like, e.g., eig in Matlab, cannot be used to compute eigenvalues of singular matrix pencils. The method we use is a projection to a regular problem of a normal rank size from [15].
4 Relation to a two-parameter eigenvalue problem
In this section we will show the main result that 2D points are eigenvalues of a related singular two-parameter eigenvalue problem. We will exploit this relation to construct numerical methods that can compute all 2D points or just a subset of ZGV points of a pencil $A+\lambda B+\mu C$.
If we assume $\mu=\mu(\lambda)$ and $x=x(\lambda)$ in (1) and differentiate $\big(A+\lambda B+\mu(\lambda)C\big)x(\lambda)=0$, we obtain (6). At a ZGV point $(\lambda,\mu)$, where $\mu^{\prime}(\lambda)=0$, we get $\big(A+\lambda B+\mu(\lambda)C\big)x^{\prime}(\lambda)+Bx(\lambda)=0$ that we write in a block form
$\left(\left[\begin{matrix}A&0\cr B&A\end{matrix}\right]+\lambda\left[\begin{matrix}B&0\cr 0&B\end{matrix}\right]+\mu\left[\begin{matrix}C&0\cr 0&C\end{matrix}\right]\right)\left[\begin{matrix}x(\lambda)\cr x^{\prime}(\lambda)\end{matrix}\right]=0.$ (14)
If we put together (1) and (14), we obtain a 2EP
$\begin{split}(A+\lambda B+\mu C)x&=0\cr\left(\left[\begin{matrix}A&0\cr B&A\end{matrix}\right]+\lambda\left[\begin{matrix}B&0\cr 0&B\end{matrix}\right]+\mu\left[\begin{matrix}C&0\cr 0&C\end{matrix}\right]\right)\left[\begin{matrix}z_{1}\cr z_{2}\end{matrix}\right]&=0.\end{split}$ (15)
Note that because of (8) all 2D points of type a), b), and c) give solutions of (15). The 2EP (15) is related to the pair of GEPs (12), where the corresponding $2n^{2}\times 2n^{2}$ matrices (13) are
$\begin{split}\Delta_{0}&=B\otimes\left[\begin{matrix}C&0\cr 0&C\end{matrix}\right]-C\otimes\left[\begin{matrix}B&0\cr 0&B\end{matrix}\right],\\ \Delta_{1}&=C\otimes\left[\begin{matrix}A&0\cr B&A\end{matrix}\right]-A\otimes\left[\begin{matrix}C&0\cr 0&C\end{matrix}\right],\\ \Delta_{2}&=A\otimes\left[\begin{matrix}B&0\cr 0&B\end{matrix}\right]-B\otimes\left[\begin{matrix}A&0\cr B&A\end{matrix}\right].\end{split}$ (16)
The GEPs $(\Delta_{1}-\lambda\Delta_{0})z=0$ and $(\Delta_{2}-\mu\Delta_{0})z=0$ are singular and the related 2EP (15) is thus singular. It however has finitely many finite eigenvalues and these include not only ZGV but also all 2D points of (1).
Theorem 4.1.
Let the $n\times n$ matrices $A,B,C$ be such that $B$ and $C$ are nonsingular and for all but a finite number of values $\lambda_{0}$ the GEP (2) has $n$ simple eigenvalues. Then the following holds for the two-parameter eigenvalue problem (15) and the related $2n^{2}\times 2n^{2}$ operator determinants (16).
  1. [left=0em]
  2. 1)
    The normal rank of the matrix pencil $\Delta_{1}-\lambda\Delta_{0}$ is $2n^{2}-n$.
  3. 2)
    If $(\lambda_{0},\mu_{0})\in\mathbb{C}^{2}$ is a 2D point for $A+\lambda B+\mu C$, then $\lambda_{0}$ is an eigenvalue of $\Delta_{1}-\lambda\Delta_{0}$.
Proof.
We can write
$\Delta_{1}-\lambda\Delta_{0}=C\otimes\left[\begin{matrix}A+\lambda B&0\cr B&A+\lambda B\end{matrix}\right]-(A+\lambda B)\otimes\left[\begin{matrix}C&0\cr 0&C\end{matrix}\right].$ (17)
For a generic $\lambda_{0}\in\mathbb{C}$ all eigenvalues of the GEP (2) are algebraically simple. Let $(A+\lambda_{0}B+\mu_{i}C)q_{i}=0$ for $i=1,\ldots,n$, where $q_{1},\ldots,q_{n}$ are nonzero eigenvectors and $\mu_{1},\ldots,\mu_{n}$ are pairwise distinct eigenvalues of the GEP (2). Then it is easy to see that
$(\Delta_{1}-\lambda_{0}\Delta_{0})\left(q_{i}\otimes\left[\begin{matrix}0\cr q_{j}\end{matrix}\right]\right)=Cq_{i}\otimes\left[\begin{matrix}0\cr(\mu_{i}-\mu_{j})Cq_{j}\end{matrix}\right]$ (18)
and
$(\Delta_{1}-\lambda_{0}\Delta_{0})\left(q_{i}\otimes\left[\begin{matrix}q_{j}\cr 0\end{matrix}\right]\right)=Cq_{i}\otimes\left[\begin{matrix}(\mu_{i}-\mu_{j})Cq_{j}\cr Bq_{j}\end{matrix}\right]$ (19)
for $i,j=1,\ldots,n$. Since $B$ and $C$ are nonsingular, vectors from (18) for $i\neq j$ and vectors from (19) are linearly independent, thus $\mathop{\rm rank}\nolimits(\Delta_{1}-\lambda_{0}\Delta_{0})\geq 2n^{2}-n$ and $\dim(\mathop{\rm Ker}\nolimits(\Delta_{1}-\lambda_{0}\Delta_{0}))\leq n$. From (18) we see that $n$ linearly independent vectors
$q_{i}\otimes\left[\begin{matrix}0\cr q_{i}\end{matrix}\right],\quad i=1,\ldots,n,$ (20)
belong to $\mathop{\rm Ker}\nolimits(\Delta_{1}-\lambda_{0}\Delta_{0})$, thus vectors (20) are basis for $\mathop{\rm Ker}\nolimits(\Delta_{1}-\lambda_{0}\Delta_{0})$ and $\mathop{\rm rank}\nolimits(\Delta_{1}-\lambda_{0}\Delta_{0})=2n^{2}-n$, which proves 1). We also see that the nullspace of the pencil $\Delta_{1}-\lambda\Delta_{0}$, see, e.g., [5], which is the set
${\cal N}=\{z(\lambda)\in\mathbb{C}(\lambda)^{2n^{2}}:\ (\Delta_{1}-\lambda\Delta_{0})z(\lambda)\equiv 0\},$
contains only vectors that have $n$ zero blocks of size $n$ accordingly to (20).
For 2), let $(\lambda_{0},\mu_{0})$ be a 2D point. We consider two options. First, if the point is of type a), b), or c), then option b) in the proof of Lemma 2.4 yields the existence of nonzero vectors $q$ and $z$ such that $(A+\lambda_{0}B+\mu_{0}C)q=0$ and $(A+\lambda_{0}B+\mu_{0}C)z+Bq=0$. It follows that $(A+\lambda_{0}B)q=-\mu_{0}Cq$ and $(A+\lambda_{0}B)z=-(Bq+\mu_{0}Cz)$. It is now easy to see that
$(\Delta_{1}-\lambda_{0}\Delta_{0})\left(q\otimes\left[\begin{matrix}q\cr z\end{matrix}\right]\right)=0$
and we have a vector in the kernel of $\Delta_{1}-\lambda_{0}\Delta_{0}$ that is clearly linearly independent from the vectors of the form (20). So, $\mathop{\rm rank}\nolimits(\Delta_{1}-\lambda_{0}\Delta_{0})<\mathop{\rm nrank}\nolimits(\Delta_{1},\Delta_{0})$ and $\lambda_{0}$ is indeed an eigenvalue of $\Delta_{1}-\lambda\Delta_{0}$.
The remaining option is that $(\lambda_{0},\mu_{0})$ is a 2D point of type d), thus the geometric multiplicity $g$ of $\mu_{0}$ as an eigenvalue of (2) is at least $2$. If $g=a$, where $a$ is the algebraic multiplicity of $\mu_{0}$, and $q_{1},\ldots,q_{g}$ are the corresponding linearly independent eigenvectors, then we have additional $g^{2}-g$ vectors
$q_{i}\otimes\left[\begin{matrix}0\cr q_{j}\end{matrix}\right]\quad{\rm for}\quad 1\leq i,j\leq g,\ i\neq j$
in the kernel of $\Delta_{1}-\lambda_{0}\Delta_{0}$ that are linearly independent from vectors (20). Therefore, $\mathop{\rm rank}\nolimits(\Delta_{1}-\lambda_{0}\Delta_{0})<\mathop{\rm nrank}\nolimits(\Delta_{1},\Delta_{0})$ and $\lambda_{0}$ is an eigenvalue. In case $a>g$ we could similarly construct enough linearly independent vectors from the kernel of $\Delta_{1}-\lambda_{0}\Delta_{0}$ to show that $\mathop{\rm rank}\nolimits(\Delta_{1}-\lambda_{0}\Delta_{0})<\mathop{\rm nrank}\nolimits(\Delta_{1},\Delta_{0})$. We omit the details and refer to [25, Thm. 4], where a similar approach was used in the proof. ∎
The assumptions in Theorem 4.1 are satisfied for generic matrices $A,B,C$, for details see Theorem A.1. In the generic case a bivariate pencil $A+\lambda B+\mu C$ has n(n1)n(n-1) 2D points. It thus follows from Theorem 4.1 that for generic matrices $A,B,C$ the normal rank of $\Delta_{1}-\lambda\Delta_{0}$ is 2n(n1)2n(n-1) and the related singular GEP $(\Delta_{1}-\lambda\Delta_{0})z=0$ has n(n1)n(n-1) finite eigenvalues. Each eigenvalue $\lambda_{0}$ of the GEP $(\Delta_{1}-\lambda\Delta_{0})z=0$ is a candidate for the $\lambda$-coordinate of a 2D point (or a ZGV point). To get the $\mu$-coordinate we fix $\lambda$ to $\lambda_{0}$ and solve the GEP (2). If $\mu_{0}$ is an eigenvalue of (2) and $x$ and $y$ are the corresponding right and left eigenvector, then:
  • [left=0em]
  • a)
    $(\lambda_{0},\mu_{0})$ is a ZGV point if $\mu_{0}$ is a simple eigenvalue and $y^{H}Bx=0$. Note that if $\mu_{0}$ is simple then directions of vectors $x$ and $y$ are unique and the test $y^{H}Bx=0$ is well defined.
  • b)
    $(\lambda_{0},\mu_{0})$ is a 2D point if the geometric multiplicity of $\mu_{0}$ is greater than one or if $y^{H}Bx=0$.
Based on the above we devised Algorithm 1 that can compute all ZGV (or 2D) points.
Algorithm 1: Compute ZGV (or 2D) points of pencil $A+\lambda B+\mu C$.
Input: $A,B,C\in\mathbb{C}^{n\times n}$, thresholds $\delta,\eta>0$.
Output: List of ZGV (or 2D) points $(\lambda_{i},\mu_{i})$, $i=1,\ldots,r$.
1: Compute $2n^{2}\times 2n^{2}$ matrices $\Delta_{0}$, $\Delta_{1}$, $\Delta_{2}$ from (16) 2: Compute eigenvalues $\lambda_{i}$, $i=1,\dots,m$, of $\Delta_{1}z=\lambda\Delta_{0}z$. 3: for $i=1,\ldots,m$: 4:    Compute eigentriples $(\mu_{j},x_{j},y_{j})$, $j=1,\ldots,n$, of $(A+\lambda_{i}B)+\mu C$ 5:    for $j=1,\ldots,n$: 6 (ZGV):     If $|y_{j}^{H}Bx_{j}|\leq\delta\|B\|_{2}$, $\sigma_{n-1}(A+\lambda_{i}B+\mu_{j}C)>\eta\|A+\lambda_{i}B+\mu_{j}\|_{2}$,      and $|y_{j}^{H}Cx_{j}|>\delta\|C\|_{2}$, then add $(\lambda_{i},\mu_{j})$ to the list. 6 (2D):     If $|y_{j}^{H}Bx_{j}|\leq\delta\|B\|_{2}$ or $\sigma_{n-1}(A+\lambda_{i}B+\mu_{j}C)\leq\eta\|A+\lambda_{i}B+\mu_{j}\|_{2}$,      then add $(\lambda_{i},\mu_{j})$ to the list.
In the following we give some additional details on Algorithm 1. In line 2 we need to compute the finite eigenvalues of a singular pencil $\Delta_{1}-\lambda\Delta_{0}$. For this purpose the projection algorithm from [15] can be applied, which we will provide as Algorithm 2 in Section 4.1.
The test for the ZGV point in line 6 is based on Lemma 2.1 and reveals if $\mu_{j}$ is a simple eigenvalue of $(A+\lambda_{i}B)+\mu C$. We detect $\mu_{j}$ as geometrically simple if the second smallest singular value of $A+\lambda_{i}B+\mu_{j}C$ is large enough.
Up to our knowledge, Algorithm 1 is the first method capable of computing all ZGV (or 2D) points of a given pencil (1). As it involves a singular GEP with the $\Delta$-matrices of size $2n^{2}\times 2n^{2}$, the method is feasible only for problems of moderate size. In the next subsection we will show that for the particular problem (15) it is possible to construct a regular 2EP whose eigenvalues include the eigenvalues of (15). This way standard numerical methods for regular 2EPs can be used to compute ZGV (2D) points.
4.1 Projected regular 2EP
In [15] a numerical method for a singular GEP using random projection of size of the normal rank is presented. If we apply it to compute the finite eigenvalues of the singular pencil $\Delta_{1}-\lambda\Delta_{0}$ of normal rank $2n^{2}-n$ related to the 2EP (15), we obtain the following algorithm.
Algorithm 2: Computing finite eigenvalues of $\Delta_{1}-\lambda\Delta_{0}$ by projection.
Input: $2n^{2}\times 2n^{2}$ matrices $\Delta_{0}$ and $\Delta_{1}$ from (16), thresholds $\delta_{1},\delta_{2}>0$.
Output: Finite eigenvalues of $\Delta_{1}-\lambda\Delta_{0}$.
1: Select random unitary $2n^{2}\times 2n^{2}$ matrices $W\ \,W_{\perp}]$ and $Z\ \,Z_{\perp}]$, where $W$ and $Z$ have $2n^{2}-n$ columns. 2: Compute the eigenvalues $\lambda_{i}$, $i=1,\dots,2n^{2}-n$, and normalized right and left eigenvectors $x_{i}$ and $y_{i}$ of $W^{H}(\Delta_{1}-\lambda\Delta_{0})Z$. 3: for $i=1,\ldots,2n^{2}-n$: 4:    $\alpha_{i}=\|W_{\perp}^{H}(\Delta_{1}-\lambda_{i}\Delta_{0})Zx_{i}\|_{2}$ and $\beta_{i}=\|y_{i}^{H}W^{H}(\Delta_{1}-\lambda_{i}\Delta_{0})Z_{\perp}\|_{2}$ 5:    $\gamma_{i}=|y_{i}^{H}W^{H}\Delta_{0}Zx_{i}|\,(1+|\lambda_{i}|^{2})^{-1/2}$ 6:    If $\max(\alpha_{i},\beta_{i})<\delta_{1}\,(\|\Delta_{1}\|_{2}+|\lambda_{i}|\,\|\Delta_{0}\|_{2})$ and $\gamma_{i}>\delta_{2}$, then add $\lambda_{i}$ to the    output list.
Algorithm 2 is based on Theorem 4.5 from [15] that shows that for generic matrices $W\ \,W_{\perp}]$ and $Z\ \,Z_{\perp}]$ the pencil $W^{H}(\Delta_{1}-\lambda\Delta_{0})Z$ is regular and contains all eigenvalues of the original singular pencil $\Delta_{1}-\lambda\Delta_{0}$. Moreover, the true finite eigenvalues of the original pencil can be separated from the true infinite eigenvalues and additional fake random eigenvalues using criteria in lines 4, 5, and 6.
In the following we will show that it suffices to take matrices $W\ \,W_{\perp}]=I\otimes[U\ \,U_{\perp}]$ and $Z\ \,Z_{\perp}]=I\otimes[V\ \,V_{\perp}]$, where $U\ \,U_{\perp}]$ and $V\ \,V_{\perp}]$ are random unitary $2n\times 2n$ matrices such that $U$ and $V$ have 2n12n-1 columns. Due to the Kronecker structure of $W$ and $Z$ we can project the initial 2EP (15) into a regular 2EP and thus use all available subspace methods for 2EPs, for instance the Jacobi-Davidson method [13, 17] and the Sylvester-Arnoldi method [22], to compute a small subset of eigenvalues close to a given target. The following theorem shows that the above choice of matrices $W$ and $Z$ leads to a regular pencil $W^{H}(\Delta_{1}-\lambda\Delta_{0})Z$.
Theorem 4.2.
For $n\times n$ matrices $A,B,C$ that satisfy the assumptions of Theorem 4.1 the following holds for the two-parameter eigenvalue problem (15) and the related $2n^{2}\times 2n^{2}$ operator determinants (16).
  1. [left=0em]
  2. 1)
    For generic $2n\times 2n$ unitary matrices $U\ \,U_{\perp}]$ and $V\ \,V_{\perp}]$, where $U$ and $V$ have 2n12n-1 columns, is the two-parameter eigenvalue problem
    $\begin{split}(A+\lambda B+\mu C)x_{1}&=0\cr\left(U^{H}\left[\begin{matrix}A&0\cr B&A\end{matrix}\right]V+\lambda\,U^{H}\left[\begin{matrix}B&0\cr 0&B\end{matrix}\right]V+\mu\,U^{H}\left[\begin{matrix}C&0\cr 0&C\end{matrix}\right]V\right)x_{2}&=0\end{split}$ (21)
    regular.
  3. 2)
    If $(\lambda_{0},\mu_{0})\in\mathbb{C}^{2}$ is an eigenvalue of (21) with a right eigenvector $x_{1}\otimes x_{2}$ and a left eigenvector $y_{1}\otimes y_{2}$, then $(\lambda_{0},\mu_{0})$ is a finite eigenvalue of (15) if and only if
    $U_{\perp}^{H}\left[\begin{matrix}A+\lambda_{0}B+\mu_{0}C&0\cr B&A+\lambda_{0}B+\mu_{0}C\end{matrix}\right]Vx_{2}=0,$ (22)
    $y_{2}^{H}U^{H}\left[\begin{matrix}A+\lambda_{0}B+\mu_{0}C&0\cr B&A+\lambda_{0}B+\mu_{0}C\end{matrix}\right]V_{\perp}=0,$ (23)
    and
    $y_{1}^{H}Bx_{1}\cdot y_{2}^{H}U^{H}\left[\begin{matrix}C&\cr&C\end{matrix}\right]Vx_{2}-y_{1}^{H}Cx_{1}\cdot y_{2}^{H}U^{H}\left[\begin{matrix}B&\cr&B\end{matrix}\right]Vx_{2}\neq 0.$ (24)
Proof.
[…]
Let us show that $(\Delta_{1}-\lambda_{0}\Delta_{0})(I\otimes V)$ has full rank. This is equivalent to
$\mathop{\rm Ker}\nolimits(\Delta_{1}-\lambda_{0}\Delta_{0})\cap\mathop{\rm Im}\nolimits(I\otimes V)=\{0\}.$ (25)
[…]
$(I\otimes V_{\perp})^{H}M=Q\mathop{\rm diag}\nolimits(v_{b}^{H}q_{1},\ldots,v_{b}^{H}q_{n}),$ (26)
where $Q=\left[\begin{matrix}q_{1}&\cdots&q_{n}\end{matrix}\right]$. Since in the generic case $v_{b}^{H}q_{i}\neq 0$ for $i=1,\ldots,n$, is the matrix (26) nonsingular and such $r$ cannot exist. Therefore, $(\Delta_{1}-\lambda_{0}\Delta_{0})(I\otimes V)$ has full rank and $(\Delta_{1}-\lambda_{0}\Delta_{0})(I\otimes V)z\neq 0$ for all $z\neq 0$.
In a similar way we can show that $\mathop{\rm Ker}\nolimits((I\otimes U)^{H})\cap\mathop{\rm Im}\nolimits(\Delta_{1}-\lambda_{0}\Delta_{0})=\{0\}$. This is equivalent to the condition $\mathop{\rm Im}\nolimits(I\otimes U)\cap\mathop{\rm Ker}\nolimits((\Delta_{1}-\lambda_{0}\Delta_{0})^{H})=\{0\}$, which is similar to (25) and we prove it in an analogous way.
For 2), we know from [15, Thm. 4.5], see also lines 3 and 5 in Algorithm 2, that
$(I\otimes U_{\perp})^{H}(\Delta_{1}-\lambda_{0}\Delta_{0})(I\otimes V)(x_{1}\otimes x_{2})=0.$ (27)
It follows from (17) and $(A+\lambda_{0}B)x_{1}=-\mu_{0}Cx_{1}$ that (27) is equal to
$Cx_{1}\otimes U_{\perp}^{H}\left[\begin{matrix}A+\lambda_{0}B+\mu_{0}C&0\cr B&A+\lambda_{0}B+\mu_{0}C\end{matrix}\right]Vx_{2}=0.$
Since $C$ is generic, $Cx_{1}\neq 0$ and (22) must hold. In a similar way, (23) follows from the condition
$(y_{1}\otimes y_{2})^{H}(I\otimes U)^{H}(\Delta_{1}-\lambda_{0}\Delta_{0})(I\otimes V_{\perp})=0.$
Finally, from [15, Thm. 4.5], see also lines 4 and 5 in Algorithm 2, it follows that
$0$ $\neq(y_{1}\otimes y_{2})^{H}(I\otimes U)^{H}\Delta_{0}(I\otimes V)(x_{1}\otimes x_{2})$
$=y_{1}^{H}Bx_{1}\cdot y_{2}^{H}U^{H}\left[\begin{matrix}C&\cr&C\end{matrix}\right]Vx_{2}-y_{1}^{H}Cx_{1}\cdot y_{2}^{H}U^{H}\left[\begin{matrix}B&\cr&B\end{matrix}\right]Vx_{2},$
which is the condition (24). ∎
This leads to the following algorithm for the computation of 2D points. If we want to compute only ZGV points, we check for each obtained 2D point $(\lambda_{0},\mu_{0})$ if $\mu_{0}$ is a simple eigenvalue of the GEP (2).
Algorithm 3: Compute 2D points of pencil $A+\lambda B+\mu C$.
Input: $A,B,C\in\mathbb{C}^{n\times n}$, thresholds $\delta_{1},\delta_{2}>0$.
Output: List of 2D points $(\lambda_{i},\mu_{i})$, $i=1,\ldots,r$.
1: Select random unitary $2n\times 2n$ matrices $U\ \,U_{\perp}]$ and $V\ \,V_{\perp}]$, where $U$ and $V$ have 2n12n-1 columns. 2: Compute the eigenvalues $(\lambda_{i},\mu_{i})$, $i=1,\dots,m$, and normalized right and left eigenvectors $x_{i1}\otimes x_{i2}$ and $y_{i1}\otimes y_{i2}$ of the 2EP (21). 3: for $i=1,\ldots,m$: 4:    $\alpha_{i}=U_{\perp}^{H}\left[\begin{matrix}A+\lambda_{i}B+\mu_{i}C&0\cr B&A+\lambda_{i}B+\mu_{i}C\end{matrix}\right]Vx_{i2}$ 5:    $\beta_{i}=y_{i2}^{H}U^{H}\left[\begin{matrix}A+\lambda_{i}B+\mu_{i}C&0\cr B&A+\lambda_{i}B+\mu_{i}C\end{matrix}\right]V_{\perp}$ 6:    $\gamma_{i}=y_{i1}^{H}Bx_{i1}\cdot y_{i2}^{H}U^{H}\left[\begin{matrix}C&\cr&C\end{matrix}\right]Vx_{i2}-y_{i1}^{H}Cx_{i1}\cdot y_{i2}^{H}U^{H}\left[\begin{matrix}B&\cr&B\end{matrix}\right]Vx_{i2}$ 7:    If ${\rm max}(\|\alpha_{i}\|_{2},\|\beta_{i}\|_{2})<\delta_{1}(\|A\|_{2}+|\lambda_{i}|\|B\|_{2}+|\mu_{i}|\|C\|_{2})$ and    $|\gamma_{i}|>\delta_{2}(1+|\lambda_{i}|^{2})^{1/2}$, then add $(\lambda_{i},\mu_{i})$ to the list of 2D points.
In line 2 of Algorithm 3 we can compute all ($m=2n^{2}-n$) or just a small subset ($m\ll 2n^{2}-n$) of the eigenvalues. The first option can be applied to problems of small size as direct methods for 2EP explicitly compute the $\Delta$-matrices and then apply the QZ algorithm. Compared to Algorithm 1, the $\Delta$-matrices in Algorithm 3 are of size $(2n^{2}-n)\times(2n^{2}-n)$, while the size in Algorithm 1 is $2n^{2}\times 2n^{2}$. The difference in size is small, but more important is that the 2EP in Algorithm 3 is nonsingular, whereas the GEP in Algorithm 1 is singular.
The nonsingularity of the 2EP in Algorithm 3 opens up options for subspace methods, see, e.g., [13, 17, 22], that compute just a subset of the eigenvalues close to a given target. These methods do not compute the $\Delta$-matrices explicitly and can thus be applied to larger problems.
5 A zero-residual Gauss-Newton method for 2D points
If we are looking for 2D points of a bivariate pencil (1), then, by Definition 2.3, we are searching for scalars $\lambda,\mu\in\mathbb{C}$ and vectors $x,y\in\mathbb{C}^{n}$ that satisfy
$\begin{split}(A+\lambda B+\mu C)x=0,\cr y^{H}(A+\lambda B+\mu C)=0,\cr y^{H}Bx=0,\cr x^{H}x=1,\cr y^{H}y=1.\cr\end{split}$ (28)
There is one equation more than the number of unknowns in $\lambda,\mu,x,y$, therefore (28) is an overdetermined nonlinear system. In addition, it is a zero-residual system since a 2D point and the corresponding right and left eigenvector solve (28) exactly. The Gauss-Newton method can be applied to compute a 2D point from an initial approximation $(\lambda_{0},\mu_{0})$ for the 2D point and $x_{0},y_{0}$ for the right and left eigenvector.
As the equations in (28) are not complex differentiable in $x$ and $y$, in order to get the Jacobian matrix, we introduce random vectors $a$ and $b$ (we can assume that they are not orthogonal to eigenvectors $x$ and $y$, respectively) to replace the normalizing conditions, define $w=\overline{y}$, and rewrite (28) as
$\begin{split}(A+\lambda B+\mu C)x=0,\cr(A^{T}+\lambda B^{T}+\mu C^{T})w=0,\cr w^{T}Bx=0,\cr a^{H}x=1,\cr b^{H}w=1.\cr\end{split}$ (29)
Suppose that we have an initial approximation $(\lambda_{0},\mu_{0},x_{0},w_{0})$ for the solution of (29). Then we get a correction $(\Delta\lambda_{0},\Delta\mu_{0},\Delta x_{0},\Delta w_{0})$ for the update
$(\lambda_{1},\mu_{1},x_{1},w_{1})=(\lambda_{0},\mu_{0},x_{0},w_{0})+(\Delta\lambda_{0},\Delta\mu_{0},\Delta x_{0},\Delta w_{0})$
from the $(2n+3)\times(2n+2)$ least squares problem
$J_{F}(\lambda_{0},\mu_{0},x_{0},w_{0})\Delta s_{0}=-F(\lambda_{0},\mu_{0},x_{0},w_{0}),$
where the Jacobian matrix is
$J_{F}(\lambda_{0},\mu_{0},x_{0},w_{0})=\left[\begin{matrix}A+\lambda_{0}B+\mu_{0}C&0&Bx_{0}&Cx_{0}\cr 0&A^{T}+\lambda_{0}B^{T}+\mu_{0}C^{T}&B^{T}w_{0}&C^{T}w_{0}\cr w_{0}^{T}B&x_{0}^{T}B^{T}&0&0\cr a^{H}&0&0&0\cr 0&b^{H}&0&0\end{matrix}\right]$ (30)
and
$\Delta s_{0}=\left[\begin{matrix}\Delta x_{0}\cr\Delta w_{0}\cr\Delta\lambda_{0}\cr\Delta\mu_{0}\end{matrix}\right],\quad F(\lambda_{0},\mu_{0},x_{0},w_{0})=\left[\begin{matrix}(A+\lambda_{0}B+\mu_{0}C)x_{0}\cr(A^{T}+\lambda_{0}B^{T}+\mu_{0}C^{T})w_{0}\cr w_{0}^{T}Bx_{0}\cr a^{H}x_{0}-1\cr b^{H}w_{0}-1\end{matrix}\right].$
The update $\Delta s_{0}$ is the least squares solution, i.e., $\Delta s_{0}=-J_{F}(\lambda_{0},\mu_{0},x_{0},w_{0})^{+}F(\lambda_{0},\mu_{0},x_{0},w_{0})$ using pseudoinverse of the Jacobian matrix. The method is given in Algorithm 4.
Algorithm 4: Compute a 2D point of pencil $A+\lambda B+\mu C$.
Input: $A,B,C\in\mathbb{C}^{n\times n}$, initial approximations $(\lambda_{0},\mu_{0})$ for the 2D point and $x_{0},y_{0}$ for the right and left eigenvector, random vectors $a,b$
Output: A 2D point $(\lambda,\mu)$
1: Set $w_{0}=\overline{y}_{0}$ 2: for $k=0,1,2,\ldots$ until convergence do 3:    Compute     $J_{F}=\left[\begin{matrix}A+\lambda_{k}B+\mu_{k}C&0&Bx_{k}&Cx_{k}\cr 0&A^{T}+\lambda_{k}B^{T}+\mu_{k}C^{T}&B^{T}w_{k}&C^{T}w_{k}\cr w_{k}^{T}B&x_{k}^{T}B^{T}&0&0\cr a^{H}&0&0&0\cr 0&b^{H}&0&0\end{matrix}\right],$ 4:    $F=\left[\begin{matrix}(A+\lambda_{k}B+\mu_{k}C)x_{k}\cr(A^{T}+\lambda_{k}B^{T}+\mu_{k}C^{T})w_{k}\cr w_{k}^{T}Bx_{k}\cr a^{H}x_{k}-1\cr b^{H}w_{k}-1\end{matrix}\right],\qquad\left[\begin{matrix}\Delta x_{k}\cr\Delta w_{k}\cr\Delta\lambda_{k}\cr\Delta\mu_{k}\end{matrix}\right]=-J_{F}^{+}F$ 5:    $x_{k+1}=x_{k}+\Delta x_{k}$, $w_{k+1}=w_{k}+\Delta w_{k}$, $\lambda_{k+1}=\lambda_{k}+\Delta\lambda_{k}$, $\mu_{k+1}=\mu_{k}+\Delta\mu_{k}$
A numerical method that provides good initial approximations $(\lambda_{0},\mu_{0})$ for 2D points is the MFRD from Section 6. Beside an approximation for a 2D point, Algorithm 4 requires initial approximations for the right and left eigenvector as well. If they are not provided, we use the following heuristics to set the initial vectors $x_{0}$ and $y_{0}$. Let $U\Sigma V^{T}$ be a singular value decomposition of $A+\lambda_{0}B+\mu_{0}C$. If $\sigma_{n-1}\ll\sigma_{n-2}$ or $\sigma_{n-1}\approx\sigma_{n}$, this suggest that a possible 2D point is of type c) or d). In this case we take for $x_{0}$ a random combination of the right singular vectors $v_{n-1}$ and $v_{n}$ such that $\|x_{0}\|_{2}=1$ and then select for $y_{0}$ a linear combination of $u_{n-1}$ and $u_{n}$ such that $y_{0}^{H}Bx_{0}=0$ and $\|y_{0}\|_{2}=1$. Otherwise, we assume that a 2D point is of type a) or b), and take $x_{0}=v_{n}$ and $y_{0}=u_{n}$.
In general the convergence of the Gauss-Newton method is not guaranteed. It is known however that the Gauss-Newton method converges locally quadratically for a zero-residual problem if the Jacobian $J_{F}$ has full rank at the solution, see, e.g., [6, Section 4.3.2] or [26, Section 10.4]. We can show that the Jacobian $J_{F}(\lambda_{0},\mu_{0},x_{0},y_{0})$ has full rank at a generic 2D point of type a), where $\lambda_{0}$ is a double eigenvalue of (3), $(\lambda_{0},\mu_{0})$ is a ZGV point and $x_{0}$ and $y_{0}$ are the corresponding right and left eigenvector. To show this, we need the following auxiliary result.
Lemma 5.1.
Let $\xi\in\mathbb{C}$ be an eigenvalue of algebraic multiplicity two and geometric multiplicity one of an $n\times n$ regular matrix pencil $A-\lambda B$. Let nonzero vectors $x,y,s,t\in\mathbb{C}^{n}$ be respectively the right and the left eigenvector and the left and the right generalized eigenvectors of order two such that
$\begin{split}(A-\xi B)x&=0,\\ (A-\xi B)s&=Bx,\\ y^{H}(A-\xi B)&=0,\\ t^{H}(A-\xi B)&=y^{H}B.\end{split}$ (31)
Then, $y^{H}Bs=t^{H}Bx\neq 0$.
Proof.
The equality $y^{H}Bs=t^{H}Bx$ follows if we multiply the second equation of (31) by $t^{H}$ from the left and the fourth equation by $s$ from the right.
To show that the quantity is nonzero, we use the Kronecker canonical form (KCF) of the pencil $A-\lambda B$, see, e.g., [9]. If follows from the KCF of $A-\lambda B$ that there exist nonsingular matrices $P$ and $Q$ such that
$P(A-\lambda B)Q=\left[\begin{matrix}J_{2}(\xi)&\cr&C-\lambda D\end{matrix}\right],\quad{\rm where}\ J_{2}(\xi)=\left[\begin{matrix}\xi-\lambda&1\cr&\xi-\lambda\end{matrix}\right]$
and $C-\lambda D$ is a regular $(n-2)\times(n-2)$ block diagonal pencil that contains the remaining Jordan and infinite blocks and such that $\xi$ is not an eigenvalue of $C-\lambda D$. It is easy to see that
$x$ $=\alpha_{1}Qe_{1},\quad s=\alpha_{1}Qe_{2}+\alpha_{2}Qe_{1},$
$y$ $=\beta_{1}P^{H}e_{2},\quad t=\beta_{1}P^{H}e_{1}+\beta_{2}P^{H}e_{2}$
for some scalars $\alpha_{1},\alpha_{2},\beta_{1},\beta_{2}$, where $\alpha_{1}$ and $\beta_{1}$ are nonzero. It then follows from the second equation of (31) that
$t^{H}Bx=t^{H}(A-\xi B)s=(\beta_{1}e_{1}+\beta_{2}e_{2})^{H}\left[\begin{matrix}0&1&\cr&0&\cr&&C-\xi D\end{matrix}\right](\alpha_{1}e_{2}+\alpha_{2}e_{1})=\overline{\beta}_{1}\alpha_{1}\neq 0.$
Lemma 5.2.
Let $(\lambda_{0},\mu_{0})\in\mathbb{C}^{2}$ be a ZGV point of a biregular bivariate pencil $A+\lambda B+\mu C$, let $x_{0}$ and $y_{0}$ be the corresponding right and left eigenvector, and $w_{0}=\overline{y}_{0}$. If the algebraic multiplicity of $\lambda_{0}$ as an eigenvalue of (3) is two, then the Jacobian (30), where $a$ and $b$ are such nonzero vectors that $a^{H}x_{0}=1$ and $b^{H}w_{0}=1$, has full rank.
Proof.
Suppose that the Jacobian $J_{F}(\lambda_{0},\mu_{0},x_{0},w_{0})$ is rank deficient. Then there exist vectors $s,t$ and scalars $\alpha,\beta$, not all being zero, such that $J_{F}(\lambda_{0},\mu_{0},x_{0},w_{0})\left[\begin{matrix}s^{T}&t^{T}&\alpha&\beta\end{matrix}\right]^{T}=0,$ i.e.,
$(A+\lambda_{0}B+\mu_{0}C)s+\alpha Bx_{0}+\beta Cx_{0}$ $=0,$ (32)
$(A^{T}+\lambda B^{T}+\mu C^{T})t+\alpha B^{T}w_{0}+\beta C^{T}w_{0}$ $=0,$ (33)
$w_{0}^{T}Bs+x_{0}^{T}B^{T}t$ $=0,$ (34)
$a^{H}s$ $=0,$ (35)
$b^{H}t$ $=0.$ (36)
If we multiply (32) from the left by $y_{0}^{H}$ then Lemma 2.2 yields that $\beta=0$ as $(\lambda_{0},\mu_{0})$ is a ZGV point.
If $\alpha\neq 0$, then it follows from (32) and (33) that $s$ and $\overline{t}$ are (up to a multiplication by a nonzero scalar) left and right generalized eigenvectors of degree two of the GEP (3) for the eigenvalue $\lambda_{0}$. But then by Lemma 5.2 $w_{0}^{T}Bs=x_{0}^{T}B^{T}t\neq 0$ and (34) does not hold. Therefore, $\alpha=0$.
Since $\alpha=\beta=0$ it follows from (32) that $s=\gamma x_{0}$ for a scalar $\gamma$ and then $s=0$ because of (35). In a similar way we get from (33) and (36) that $t=0$. This shows that there are no nonzero vectors in the kernel of $J_{F}(\lambda_{0},\mu_{0},x_{0},w_{0})$. ∎
A generic ZGV point $(\lambda_{0},\mu_{0})$ is such that $\lambda_{0}$ is a double eigenvalue of (3). It follows from Lemma 5.1 that near such points Algorithm 4 converges quadratically. There might also exist ZGV points where multiplicity of $\lambda_{0}$ is higher than two. At such points we can expect a linear convergence.
Although we are primarily interested in ZGV points, Algorithm 4 converges to all types of 2D points for a bivariate pencil $A+\lambda B+\mu C$. If a 2D point is not a ZGV point, i.e., it is a 2D point of type b), c) or d), then the Jacobian (30) is rank deficient and we can expect a linear convergence.
6 Method of fixed relative distance
A method from [18], designed to compute approximations for values $\mu$ such that $A+\mu B$ has a multiple eigenvalue, can be straightforward generalized to compute good approximations for 2D points of a biregular pencil (1), as already discussed in [18, Sec. 6].
By Lemma 2.4 $(\lambda_{0},\mu_{0})\in\mathbb{C}^{2}$ is a 2D point of (1) if and only if $\lambda_{0}$ is a multiple (generically double) eigenvalue of (3). Thus, for $\widetilde{\mu}_{0}\approx\mu_{0}$ such that $\widetilde{\mu}_{0}\neq\mu_{0}$ the GEP
$\left((A+\widetilde{\mu}_{0}C)+\lambda B\right)x=0$ (37)
has two different eigenvalues $\lambda_{01}$ and $\lambda_{02}$ close to $\lambda_{0}$. In the method of fixed relative distance (MFRD) we select a regularization parameter $\delta>0$ and search for $\widetilde{\mu}_{0}$ such that (37) has eigenvalues $\lambda_{01}$ and $\lambda_{02}=(1+\delta)\lambda_{01}$. We write this as a 2EP
$\begin{split}(A+\lambda B+\mu C)x_{1}&=0,\\ (A+\lambda(1+\delta)B+\mu C)x_{2}&=0.\end{split}$ (38)
The 2EP (38) is generically nonsingular for $\delta>0$ and eigenvalues are approximations of 2D points of (1). Note that (38) has $n^{2}$ eigenvalues while generically there are only n(n1)n(n-1) 2D points of (1), so at least $n$ of the eigenvalues of (38) are not related to 2D points of (1). It is easy to see that (38) has $n$ solutions of the form $(0,\mu)$, where $\mu$ is an eigenvalue of the GEP $(A+\mu C)x=0$, which are generically not related to nearby 2D points. For each eigenvalue $(\lambda,\mu)$ of (38) we thus have to check if it corresponds to a 2D point of (1). For this task and also to refine the solution, which is inaccurate due to the regularization parameter $\delta$, we use Algorithm 4. A parallel approach is applied in [18], where approximations obtained from a regularized 2EP are refined using a similar zero-residual Gauss-Newton method.
The role of the regularization parameter $\delta$ is the following. If $\delta$ is too small, then the GEP (38) is close to being singular which makes it difficult to find eigenvalues using methods for nonsingular 2EPs. On the other hand, by increasing the value of $\delta$ approximations for the 2D points are becoming less accurate and Algorithm 4 might fail to converge to a 2D point.
7 Numerical examples and applications
We provide several numerical examples together with some applications and related problems. All numerical experiments were carried out in Matlab 2023a on a desktop PC with 64 GB RAM and i7–11700K 3.6 GHz CPU. Methods from MultiParEig [28] were used to solve the related 2EPs and singular GEPs. Some results were computed in higher precision using the Advanpix Multiprecision Computing Toolbox [1]. The code and data for all numerical examples in this paper are available at https://github.com/borplestenjak/ZGV_Points.
Example 7.1.
We start with a simple example, where we consider
$A=\left[\begin{matrix}3&0\cr 0&0\end{matrix}\right],\quad B=\left[\begin{matrix}0&1\cr-1&-1\end{matrix}\right],\quad C=\left[\begin{matrix}-2&-2\cr 2&0\end{matrix}\right].\quad$
Then $\det(A+\lambda B+\mu C)=\lambda^{2}-2\lambda\mu+4\mu^{2}-3\lambda$ and real eigenvalues $(\lambda,\mu)$ lie on the ellipse shown in Figure 2. The two ZGV points, marked on the figure, are $Z_{1}=(1,0.5)$ and $Z_{2}=(3,1.5)$.
Figure 2: Real eigenvalues and ZGV points of Example 7.1.
If we apply Algorithm 1, we get matrices $\Delta_{0},\Delta_{1},\Delta_{2}$ (16) of size $8\times 8$. The singular pencil $\Delta_{1}-\lambda\Delta_{0}$ has two finite eigenvalues $\lambda_{1}=1$ and $\lambda_{2}=3$, which we compute using Algorithm 2. By computing the corresponding eigenvalues $\mu$ of the pencils $(A+\lambda_{i}B)+\mu C$ we get both ZGV points. The maximal error $\|Z_{i}-(\lambda_{i},\mu_{i})\|_{2}$ for $i=1,2$ of the computed ZGV points is $1.6\cdot 10^{-15}$.
Alternatively, we could use Algorithm 3. The eigenvalues $(\lambda_{i},\mu_{i})$ of the projected regular 2EP are listed in Table 1 together with the values $\widetilde{\alpha}_{j}=\alpha_{j}/(\|A\|_{2}+|\lambda_{j}|\|B\|_{2}+|\mu_{j}|\|C\|_{2})$, $\widetilde{\beta}_{j}=\beta_{j}/(\|A\|_{2}+|\lambda_{j}|\|B\|_{2}+|\mu_{j}|\|C\|_{2})$, and $\widetilde{\gamma}_{j}=\gamma_{j}(1+|\lambda_{j}|^{2})^{-1/2}$ that are used to identify the two regular eigenvalues corresponding to ZGV points. The maximal error in this case is $3.0\cdot 10^{-15}$.
Table 1: Results of Algorithm 3 applied to Example 7.1.
$j$ $\lambda_{j}$ $\mu_{j}$ $\quad\ \widetilde{\alpha}_{j}$ $\quad\ \widetilde{\beta}_{j}$ $\quad\ \widetilde{\gamma}_{j}$
1 $1.00000-1.9\cdot 10^{-15}i$ $0.50000-1.2\cdot 10^{-15}i$ $3.9\cdot 10^{-16}$ $1.6\cdot 10^{-16}$ $7.9\cdot 10^{-2}$
2 $3.00000-2.3\cdot 10^{-15}i$ $\phantom{-}1.50000-6.1\cdot 10^{-16}i$ $8.9\cdot 10^{-17}$ $2.5\cdot 10^{-16}$ $3.3\cdot 10^{-1}$
3 $1.4\cdot 10^{15}-1.1\cdot 10^{15}i$ $\phantom{-}8.3\cdot 10^{14}+3.1\cdot 10^{14}i$ $7.7\cdot 10^{-16}$ $3.8\cdot 10^{-16}$ $1.3\cdot 10^{-31}$
4 $2.9\cdot 10^{15}+2.9\cdot 10^{15}i$ $\phantom{-}2.0\cdot 10^{15}-5.3\cdot 10^{14}i$ $1.2\cdot 10^{-16}$ $6.6\cdot 10^{-17}$ $3.0\cdot 10^{-32}$
5 $0.96018+0.34932i$ 0.520330.00224i-0.52033-0.00224i $4.1\cdot 10^{-2}$ $3.8\cdot 10^{-16}$ $8.2\cdot 10^{-2}$
6 4.180910.52938i4.18091-0.52938i $\phantom{-}1.46522-0.38308i$ $5.8\cdot 10^{-16}$ $7.3\cdot 10^{-2}$ $1.2\cdot 10^{-1}$
Yet another option is to apply the MFRD and Algorithm 4. If we take $\delta=10^{-2}$ and solve the 2EP (38), we get four candidates for ZGV points: $(\widetilde{\lambda}_{1},\widetilde{\mu}_{1})=(0.99503,-0.49999)$, $(\widetilde{\lambda}_{2},\widetilde{\mu}_{2})=(2.98504,1.49996)$, $(\widetilde{\lambda}_{3},\widetilde{\mu}_{3})=(-1.24\cdot 10^{-22},1.63\cdot 10^{-11})$, $(\widetilde{\lambda}_{4},\widetilde{\mu}_{4})=(0,0)$. We use them as initial values for the Gauss-Newton method in Algorithm 4 and select initial approximations for left and right eigenvectors from the SVD of $A+\widetilde{\lambda}_{i}B+\widetilde{\mu}_{i}C$. The first two candidates converge quadratically to $Z_{1}$ and $Z_{2}$, respectively, with the maximal error $1.2\cdot 10^{-16}$.
Example 7.2.
To illustrate the quadratic convergence of Algorithm 4 near a ZGV point, we apply the algorithm to the pencil (9) from Example 2.8 and initial approximations obtained by the MFRD using $\delta=10^{-2}$. The convergence plots of nine obtained solutions, composed of six ZGV points and three 2D points of type d), where the eigencurves intersect, are presented in Figure 3. To emphasize the quadratic convergence, we used computation in higher precision. Although the Jacobian is singular at the solution, we observe initial quadratic convergence for the intersection points as well, followed by steps where the error suddenly increases close to the solution. This happens because for these eigenvalues the right and left eigenvectors are not unique.
Figure 3: Convergence plot for Algorithm 4 applied to 2D points of Example 2.8.
Example 7.3.
We take random real $n\times n$ matrices $A$, $B$, $C$ and compare the timings and errors of the computed ZGV points. The exact ZGV points to compare with were computed by an additional refinement using Algorithm 4 in quadruple precision. For each $n=5,10,\ldots,45$ we solved 10 random problems. The comparison of median computational time and median maximal relative error of the obtained ZGV points is presented in Figure 4. We applied Algorithm 4 to refine the results obtained by Algorithm 1 and Algorithm 3 as well.
Figure 4: Computational times (left hand side) and errors (right hand side) on random pencils.
All three methods have complexity ${\cal O}(n^{6})$. Algorithm 1 is the most expensive since it requires the solution of a singular GEP of size $2n^{2}\times 2n^{2}$ to compute the candidates for $\lambda$. In addition, for each candidate we have to solve an $n\times n$ GEP to obtain the possible $\mu$ parts. Algorithm 3 is more efficient as we have to solve a regular and slightly smaller GEP of size $(2n^{2}-n)\times(2n^{2}-n)$, and no additional smaller GEPs have to be solved. The MFRD is asymptotically the fastest approach as the main task is to solve a regular MEP with matrices of size $n\times n$, which is equivalent to solving a regular GEP of size $n^{2}\times n^{2}$.
As expected, the MFRD combined with Algorithm 4 is the fastest method for large $n$. It is also the most accurate method due to the final refinement by the Gauss-Newton method, but note that this refinement can be applied with virtually no extra cost to Algorithm 1 and Algorithm 3 as well which reduces the error substantially. Based on these results we see that Algorithms 1 and 3 are suitable for small problems where $n\leq 10$, while for larger problems one should use the MFRD.
7.1 2D-eigenvalue problem
In [21] a 2D-eigenvalue problem (2DEVP) is studied, which turns out to be a special case of 2D points. In 2DEVP we have Hermitian matrices $A,B\in\mathbb{C}^{n\times n}$, where $B$ is indefinite, and we are searching for a pair $(\lambda,\mu)\in\mathbb{R}^{2}$, called a 2D-eigenvalue, and nonzero $x\in\mathbb{C}^{n}$ such that
$\begin{split}(A-\lambda B)x&=\mu x,\cr x^{H}Bx&=0,\cr x^{H}x&=1.\end{split}$ (39)
We can see (39) as a special case of (28). Namely, for $\lambda\in\mathbb{R}$ is the matrix $A-\lambda B$ Hermitian and has real eigenvalues $\mu_{1},\ldots,\mu_{n}$. The corresponding left and right eigenvectors are equal. Thus, each 2D-eigenvalue of (39) is a 2D point of the bivariate pencil $A+\lambda(-B)+\mu(-I)$ and each solution $(\lambda,\mu,x)$ of (39) gives a solution $(\lambda,\mu,x,x)$ of the corresponding (28). The related problem (28) can have additional solutions as $\lambda$ and $\mu$ in (28) can be complex and right and left eigenvectors $x$ and $y$ are not necessary colinear. Therefore, we can apply Algorithm 1 or Algorithm 3 to compute all solutions of (28) and then a subset of real pairs $(\lambda,\mu)$ is the solution of (39). Up to our knowledge, this is the first global method that computes all 2D-eigenvalues of a 2DEVP. All other algorithms, including the one introduced in [21], are local methods that require good initial approximations and compute a single solution of (39).
Alternatively, we can apply the MFRD from Section 6 to obtain good approximations for the solutions of a related problem (28) and then refine the approximations by applying the Gauss-Newton type method to (39) in a similar way as in Algorithm 4.
Example 7.4.
The matrices of a 2D-eigenvalue problem from [21, Ex. 1] are
$A=\left[\begin{matrix}2&0&1\cr 0&0&1\cr 1&1&0\end{matrix}\right],\quad B=\left[\begin{matrix}1&0&1\cr 0&1&1\cr 1&1&0\end{matrix}\right].\quad$
Figure 5 shows the real eigencurves of the corresponding bivariate pencil $A+\lambda(-B)+\mu(-I)$ together with the three 2D-eigenvalues, which are all 2D points of type a). The 2D point $(1,0)$ is such that $\lambda=1$ is an eigenvalue of multiplicity three for the pencil $A-\lambda B$.
Figure 5: 2D-eigenvalues of Example 7.4
[…]
Example 7.5.
We consider 2D-eigenvalues of $n\times n$ banded Toeplitz matrices
$A={\rm pentadiag}(1,0,5,0,1),\quad B={\rm tridiag}(1,1/2,1).$ (40)
For each $\lambda\in\mathbb{R}$ the eigenvalues of $A-\lambda B$ can be ordered as $\mu_{n}(\lambda)\leq\cdots\leq\mu_{1}(\lambda)$. The real eigencurves of $A-\lambda B-\mu I$ have a particular structure that is visible in Figure 6 for $n=10$ (left hand figure) and $n=20$ (close up, right hand figure). For $n=2m$ the curves $\mu_{2k-1}(\lambda)$ and $\mu_{2k}(\lambda)$ for $k=1,\ldots,m$ touch at 2k12k-1 points, which are 2D points of type d) and are represented by red dots. Together there are $m^{2}$ such 2D-eigenvalues. The additional 2D-eigenvalues are ZGV points represented by white dots, where $\mu^{\prime}(\lambda)=0$. The 2D-eigenvalues in Figure 6 were computed with the MFRD combined with Algorithm 4.
Figure 6: 2D-eigenvalues of (40) for $n=10$ (left hand side) and $n=20$ (right hand side), where white and red points respectively represent ZGV points and 2D points of type d).
For a large $n$ it is not possible to compute all 2D-eigenvalues due to the size of the $\Delta$ matrices of the corresponding 2EP. For such problems we can apply the MFRD and solve the 2EP (38) by a subspace method that computes a small number of eigenvalues close to the target. We demostrate this approach on (40) and $n=100$ (this leads to $\Delta$ matrices of size $10000\times 10000$ is we apply the MFRD, or $20000\times 20000$ if we apply Algorithm 1). The computed 2D-eigenvalues are presented in Figure 7. The left hand side figure shows the first 100 2D-eigenvalues obtained by the Krylov-Schur method for the 2EP [22] that we applied to search for the 2D-eigenvalues $(\lambda,\mu)$ with a target $\lambda_{0}=-0.15$. As it can be seen from the figure, we indeed get 2D-eigenvalues such that $\lambda$ is close to $\lambda_{0}$. The right hand side figure shows the first eight 2D-eigenvalues obtained by the Jacobi-Davidson method with harmonic Ritz values for the 2EP [17] using the target $(\lambda_{0},\mu_{0})=(0,5)$. For more details and the choice of the parameters, see the Matlab code in the repository with the numerical examples.
The computed four 2D-eigenvalues from the right hand side of Figure 7 that are closest to $(0,5)$ are:
$(\lambda_{1},\mu_{1})$ $=(-5.3655644583\cdot 10^{-2},4.99203880712),$
$(\lambda_{2},\mu_{2})$ $=(\phantom{-}5.5000908314\cdot 10^{-2},5.00489644932),$
$(\lambda_{3},\mu_{3})$ $=(-3.3513752079\cdot 10^{-2},5.04903430286),$
$(\lambda_{4},\mu_{4})$ $=(\phantom{-}3.2092945395\cdot 10^{-2},4.94904259728).$
Figure 7: 2D-eigenvalues of (40) for $n=100$. On the left hand side are first 100 solutions obtained with the Krylov-Schur method and the target $\lambda_{0}=-0.15$, and on the right hand side are first 8 solutions obtained with the Jacobi-Davidson method and the target $(\lambda_{0},\mu_{0})=(0,5)$
To access the influence of the parameter $\delta$ in the MFRD, we applied the method to (40) with $n=10$ using values $\delta=10^{-k}$ for $k=1,\ldots,10$. The problem has 39 ZGV points and 25 additional 2D points of type d). For each $\delta$ we performed 10 runs (the results vary due to random processes in the MEP solver from MultiParEig and the random vectors $a$ and $b$ in the Gauss-Newton method) and computed the empirical success rate, defined as the average ratio of recovered points (for all 2D points, 2D points of type d), and ZGV points, respectively). The results are summarized in Table 2.
We observe that $\delta$ should not be too small nor too large. If $\delta$ is too small, the MEP is nearly singular and the computed eigenvalues are inaccurate. If $\delta$ is too large the eigenvalues may be perturbed too far away for the Gauss-Newton method to converge. The success rate of recovering 2D points of type d) is higher for smaller $\delta$ because the corresponding eigenvalues have higher multiplicities and thus perturb more readily. On the other hand, the eigenvalues corresponding to ZGV points are simple and easier to detect when $\delta$ is larger.
Table 2: Empirical success rate of recovering 2D and ZGV points of the pencil $A-\lambda B-\mu I$ for matrices (40) with $n=10$ using the MFRD and Gauss-Newton.
$\delta$ $P_{\rm all}(\delta)$ $P_{\rm 2D(d)}(\delta)$ $P_{\rm ZGV}(\delta)$ $\delta$ $P_{\rm all}(\delta)$ $P_{\rm 2D(d)}(\delta)$ $P_{\rm ZGV}(\delta)$
$10^{-1}$ $0.5852$ $0.2320$ $0.8115$ $10^{-6}$ $1.0000$ $1.0000$ $1.0000$
$10^{-2}$ $0.8070$ $0.5060$ $1.0000$ $10^{-7}$ $0.9891$ $1.0000$ $0.9821$
$10^{-3}$ $0.9273$ $0.8140$ $1.0000$ $10^{-8}$ $0.7797$ $1.0000$ $0.6385$
$10^{-4}$ $1.0000$ $1.0000$ $1.0000$ $10^{-9}$ $0.5281$ $1.0000$ $0.2256$
$10^{-5}$ $1.0000$ $1.0000$ $1.0000$ $10^{-10}$ $0.5398$ $0.9800$ $0.2577$
7.2 Distance to instability
The 2DEVP is related to the problem of computing the distance to instability of a stable matrix. If $A\in\mathbb{C}^{n\times n}$ is a stable matrix, i.e., $\mathop{\rm Re}\nolimits({\lambda})<0$ for all eigenvalues $\lambda$ of $A$, then
$\beta(A)=\min\left\{\|E\|_{2}:\ A+E\textrm{ is unstable}\right\}.$
It is well known, see, e.g., [34], that $\beta(A)=\min_{\lambda\in\mathbb{R}}(\sigma_{\rm min}(A-\lambda iI)),$ which is equivalent to
$),\beta(A)=\min_{\lambda\in\mathbb{R}}(\lambda_{\rm sp}(\widetilde{A}-\lambda\widetilde{B}$)), (41)
where $\lambda_{\rm sp}$ denotes the smallest positive eigenvalue,
$\widetilde{A}=\left[\begin{matrix}0&A\cr A^{H}&0\end{matrix}\right]\quad\textrm{and}\quad\widetilde{B}=\left[\begin{matrix}0&iI\cr-iI&0\end{matrix}\right].$ (42)
[…]
Example 7.6.
We consider the matrix
$A=\left[\begin{matrix}-0.4+6i&1&&\cr 1&-0.1+i&1&\cr 0&1&-1-3i&1\cr&&1&-5+i\end{matrix}\right]$
from Example 5 in [8]. The real eigencurves of the pencil $\widetilde{A}-\lambda\widetilde{B}-\mu I$ (42) are presented in Figure 8 together with the real 2D points computed with Algorithm 1. As we see on two closeups on the right hand side, on the contrary to Example 7.5 the eigencurves do not intersect, they just come very close and veer apart. This is a well-known behaviour expected in a general case, for more details, see, e.g., [33].
Figure 8: 2D-eigenvalues of Example 7.4
The method computes all 2D points, which appear in pairs $(\lambda,\pm\mu)$. From the 2D point $(0.95301472,0.03188701)$ with the smallest positive $\mu$ (the blue star on Figure 8) we get by further refinement by Algorithm 4 that $\beta(A)=3.188701430320041\cdot 10^{-2}$, which agrees to the result in [8].
Let us remark that this does not lead to an efficient algorithm for the computation of $\beta(A)$. In theory we could apply Algorithm 3 and use Sylvester-Arnoldi or Krylov-Schur method from [22] in line 2 to search for the real 2D point $(\lambda_{0},\mu_{0})$ with the smallest $|\mu_{0}|$, but in practice this is not an easy task. As it can be seen also from Figure 7.6, we need to compute an interior eigenvalue and, in addition, there can be many complex 2D points with a very small $|\mu_{0}|$.
7.3 Double eigenvalue problem
In the double eigenvalue problem, see, e.g., [18, 25], we have matrices $A,B\in\mathbb{C}^{n\times n}$ and we are looking for values $\mu$ such that $A+\mu B$ has a multiple eigenvalue (generically, such eigenvalues have multiplicity two). It follows that points $(\lambda_{0},\mu_{0})$, where $\lambda_{0}$ is a multiple eigenvalue of $A+\mu_{0}B$, are exactly 2D points of the pencil $A+\lambda I+\mu B$.
Based on the above, we can compute values $\mu_{0}$ such that $A+\mu B$ has a multiple eigenvalue from the following singular 2EP
$\begin{split}(A+\lambda I+\mu B)x&=0\\ \left(\left[\begin{matrix}A&0\cr I&A\end{matrix}\right]+\lambda\left[\begin{matrix}I&0\cr 0&I\end{matrix}\right]+\mu\left[\begin{matrix}B&0\cr 0&B\end{matrix}\right]\right)\left[\begin{matrix}z_{1}\cr z_{2}\end{matrix}\right]&=0.\end{split}$ (43)
The above 2EP for the double eigenvalue problem can be derived like in [15, Example 7.4] from the property that if $\lambda_{0}$ is a multiple eigenvalue of $A+\mu_{0}B$ that has algebraic multiplicity one, then there exist nonzero vectors $x$ (an eigenvector) and $y$ (a generalized eigenvector of degree two) such that $(A+\mu_{0}B+\lambda_{0}I)x=0$ and $(A+\mu_{0}B+\lambda_{0}I)y+x=0$.
A similar approach to solve the double eigenvalue problem was used in [25], where a singular 2EP with equations $(A+\lambda I+\mu B)x=0$ and $(A+\lambda I+\mu B)^{2}z=0$ was applied. The second equation is linearized into $(F+\lambda G+\mu H)z=0$, where matrices $F,G,H$ are of size $3n\times 3n$, therefore, the corresponding singular GEPs with $\Delta$-matrices in [25] are of size $3n^{2}\times 3n^{2}$. The approach from [15], which applies (43), leads to singular GEPs with $\Delta$-matrices of size $2n^{2}\times 2n^{2}$ and is more efficient due to smaller matrices. Here we give an even more efficient approach. The key is to apply Algorithm 4, where we have to solve nonsingular GEPs with $\Delta$-matrices of size $(2n^{2}-n)\times(2n^{2}-n)$. While this is still less efficient than the MFRD approach [18], where we have to solve a nonsingular 2EPs with corresponding $\Delta$-matrices of size $n^{2}\times n^{2}$, it theoretically leads to exact results in exact computation.
Example 7.7.
In Matlab we use rng(3), A=randn(25), B=randn(25) to construct two random matrices. There are 600 values $\mu$ such that $A+\mu B$ has a double eigenvalue and these are the $\mu$ coordinates of the 2D points $(\lambda,\mu)$ of the pencil $A+\lambda I+\mu B$. We can compute all the points using any of the proposed algorithms, the fastest method is the MFRD combined with the Gauss-Newton method, which calculates all points in $1.52$ seconds. In Figure 9 we present all such values $\mu$ inside $0.5,0.5]\times[-0.5,0.5]$ that we computed by Algorithm 3. For the second computation we used a Krylov-Schur for 2EP in line 2 of Algorithm 3 to compute six solutions closest to the origin that are encircled by red circles. This computation takes $0.19$ seconds and returns the following six solutions $\mu$ with the minimal absolute value:
$\mu_{1,2}$ $=-5.1825645307\cdot 10^{-2}\pm 2.1720742563\cdot 10^{-2}i,$
$\mu_{3}$ $=\phantom{-}6.9303957823\cdot 10^{-2},\quad\mu_{4}=-7.5529261749\cdot 10^{-2},$
$\mu_{5,6}$ $=\phantom{-}1.1251034552\cdot 10^{-2}\pm 1.2727979158\cdot 10^{-1}i.$
This shows that in Algorithm 3 we can apply a subspace method and compute just a small number of solutions. We can do this also if we apply the MFRD, as we have shown in Example 7.5, while in Algorithm 1 this is not possible because we have a singular eigenvalue problem.
Figure 9: Points $\mu$ in the complex plane such that $A+\mu B$ from Example 7.7 has a double eigenvalue
7.4 Critical points of two-parameter Sturm-Liouville eigencurves
We consider a two-parameter Sturm-Liouville problem [3, 4] of the form
$(p(x)y^{\prime})^{\prime}+q(x)y=(\lambda r(x)+\mu)y,\quad a\leq x\leq b,$ (44)
with boundary conditions
$\cos(\alpha)y(a)-\sin(\alpha)p(a)y^{\prime}(a)=0,\quad\cos(\beta)y(b)-\sin(\beta)p(b)y^{\prime}(b)=0,$ (45)
where $p$ is continuously differentiable and positive on $[a,b]$, $q$ and $r$ are piecewise continuous on $[a,b]$, and $\alpha,\beta$ are real. The set of $(\lambda,\mu)\in\mathbb{R}^{2}$, for which there exist a nontrivial solution $y$ that satisfies (44) and (45), forms eigencurves that are a countable union of graphs of analytic functions, for details see, e.g., [4].
For a fixed $\lambda\in\mathbb{R}$ we get a regular one-parameter Sturm-Liouville problem (44), (45). It is well known that for each $n\in{\mathbb{N}}$ there exists exactly one eigenvalue $\mu_{n}(\lambda)$ such that the corresponding eigenfunction $y$ has exactly n1n-1 zeros in $(a,b)$. We also know that $\mu_{j}(\lambda)<\mu_{j+1}(\lambda)$ for each $j\in{\mathbb{N}}$. It turns out [4, Thm. 2.1] that eigencurves $\mu_{n}(\lambda)$ for $n\in{\mathbb{N}}$ are analytic functions. We are interested in critical points, where $\mu_{n}^{\prime}(\lambda)=0$, which are clearly generalizations of ZGV points. Thus, we can discretize (44), (45) into a two-parameter pencil $A+\lambda B+\mu I$ and then compute the corresponding ZGV points.
Example 7.8.
We consider the modified Mathieu equation
$y^{\prime\prime}(x)-2\lambda\cos(2x)y(x)+\mu y(x)=0$ (46)
with boundary conditions $y^{\prime}(0)=y^{\prime}(\pi/2)=0$. The first five dispersion curves $\mu_{j}(\lambda)$ for $j=1,\ldots,5$ together with real ZGV points are presented in Figure 10. Beside the trivial ZGV points of the form $(0,4k^{2})$ for $k=0,\ldots,4$ the remaining ZGV points on the figure are $(\pm 11.14606106,17.41358458)$, $(\pm 31.48781869,42.39762508)$, and $(\pm 60.12377598,78.78937721)$.
To obtain the above ZGV points, we discretized (46) by the spectral collocation on $n=25$ points using bde2mep function in MultiParEig [28], see [10] for details, and applied Algorithm 1. We refined the solutions using finer discretizations on $n=50$ and $n=100$ points, where each time a ZGV point on a coarser grid was used as an initial approximation for Algorithm 4.
Figure 10: ZGV points for the boundary value problem (46)
7.5 ZGV points for quadratic eigenvalue problems
The expression ZGV originates from engineering applications. In the study of anisotropic elastic waveguides, see, e.g., [29], we consider an eigenvalue problem
$\big(\lambda^{2}L_{2}+\lambda L_{1}+L_{0}+\omega^{2}M\big)\,u=0,$ (47)
where $L_{0}$, $L_{1}$, $L_{2}$, $M$ are Hermitian $n\times n$ matrices such that $L_{2}$ and $M$ are nonsingular, obtained by a discretization of a boundary value problem. The solution are dispersion curves $\omega=\omega(\lambda)$. Of particular interest are the zero-group-velocity (ZGV) points, where $\omega$ and $\lambda$ are real, and $\omega^{\prime}(\lambda)=0$. If we assume that $u=u(\lambda)$ and $\omega(\lambda)$ are differentiable, we obtain by differentiating (47) that at a ZGV point $(\lambda,\omega)$ it holds
$\big(\lambda^{2}\,\widetilde{L}_{2}+\lambda\,\widetilde{L}_{1}+\widetilde{L}_{0}+\omega^{2}\,\widetilde{M}\big)\,\widetilde{u}=0,$ (48)
where
$\widetilde{L}_{2}=\left[\begin{matrix}L_{2}&0\\ 0&L_{2}\end{matrix}\right],\quad\widetilde{L}_{1}=\left[\begin{matrix}L_{1}&0\\ 2L_{2}&L_{1}\end{matrix}\right],\quad\widetilde{L}_{0}=\left[\begin{matrix}L_{0}&0\\ L_{1}&L_{0}\end{matrix}\right],\quad\widetilde{M}=\left[\begin{matrix}M&0\\ 0&M\end{matrix}\right],\quad\widetilde{u}=\left[\begin{matrix}u\\ u^{\prime}\end{matrix}\right].$
Equations (47) and (48) form a quadratic 2EP [24, 16] and numerical methods similar to the presented in this paper can be derived for the computation of ZGV points of (47). For more details, see [19], where a method for the above quadratic 2EP is presented together with a generalization of the MRFD from Section 6 and a generalized Gauss–Newton method from Section 5.
Although the next approach is less efficient than the one in [19], which exploits the connection to a singular quadratic 2EP, we can substitute $\mu=\omega^{2}$ and apply any of the standard linearizations for the quadratic eigenvalue problem to transform (47) into a problem of form (1) with $2n\times 2n$ matrices. We see from $\mu^{\prime}(\lambda)=2\omega(\lambda)\omega^{\prime}(\lambda)$ that each ZGV point of (47) corresponds to a ZGV point of the linearized pencil, which can in addition have ZGV points with $\mu=0$. For instance, we can write (47) as
$\left(\left[\begin{matrix}L_{0}&L_{1}\cr 0&-I\end{matrix}\right]+\lambda\left[\begin{matrix}0&-L_{2}\cr I&0\end{matrix}\right]+\mu\left[\begin{matrix}M&0\cr 0&0\end{matrix}\right]\right)\left[\begin{matrix}u\cr\lambda u\end{matrix}\right]=0$ (49)
and apply Algorithm 1 or Algorithm 4 to (49). We illustrate this with a small example.
Example 7.9.
The problem (47) with the following diagonal and tridiagonal matrices
$L_{2}=\left[\begin{matrix}-1&0.5&\cr 0.5&-2&0.5\cr&0.5&-3\end{matrix}\right],\ L_{1}=\left[\begin{matrix}1&-0.25&\cr-0.25&2&-0.25\cr&-0.25&-3\end{matrix}\right],\ L_{0}=\left[\begin{matrix}-1&&\cr&-2&\cr&&-3\end{matrix}\right],\ M=\left[\begin{matrix}2&1&\cr 1&3&1\cr&1&4\end{matrix}\right]$}
has five real ZGV points $(\lambda,\omega)$ such that $\omega>0$, which are presented in Figure 11. By applying Algorithm 1 to the linearization (49) we compute these ZGV points as:
$(\lambda_{1},\omega_{1})$ $=(-0.2312197373,0.79089022421),\ (\lambda_{2},\omega_{2})=(0.3684223373,0.82195756940),$
$(\lambda_{3},\omega_{3})$ $=(\phantom{-}0.6315720581,0.54233673936),\ (\lambda_{4},\omega_{4})=(0.1584790129,0.82797266404),$
$(\lambda_{5},\omega_{5})$ $=(\phantom{-}0.1200999663,1.10785496051).$
Figure 11: Real ZGV points $(\lambda,\omega)$ with positive $\omega$ for the parameter-dependent quadratic eigenvalue problem from Example 7.9
8 Conclusions
We have investigated critical points of eigencurves of $n\times n$ bivariate matrix pencils and provided common theory that links ZGV points, 2D-eigenvalues from [21], and the newly introduced 2D points. We derived a singular 2EP, which is challenging to solve, whose solutions represent the critical points.
We proposed three numerical methods for the computation of critical points. The first method in Algorithm 1 computes all critical points from eigenvalues of a singular GEP of size $2n^{2}\times 2n^{2}$ that is built from operator determinants related to the singular 2EP. Eigenvalues of this singular GEP are computed via a random projection to a GEP of the size of the normal rank from [15]. Because of the high complexity it is in practice feasible only for small problems. The second method in Algorithm 3 follows the same idea, but due to a structured random projection requires the above singular GEP only implicitly. Furthermore, by projecting it into a regular 2EP, this approach is more efficient than Algorithm 1 for the computation of all critical points, and, unlike Algorithm 1, it can be used to compute just a small number of solutions close to a target. Both of the above algorithms can find all critical points in exact computation.
The third approach is a locally convergent Gauss-Newton-type method, which exhibits quadratic convergence near ZGV points. Combining this method with the method of fixed relative distance from [18] that provides initial approximations yields a solver faster than the above, which is suitable for larger problems and is likely to find all critical points. The Gauss-Newton-type method can as well be applied to refine the solutions obtained by Algorithm 1 and Algorithm 3.
We presented many possible applications. Through extensive numerical experiments, we have demonstrated that the proposed numerical methods can successfully compute either all critical points or a subset of critical points close to a given target.
Acknowledgments The author would like to thank the anonymous referees for their helpful comments and suggestions.
Funding Funding was provided by Javna agencija za znanstvenoraziskovalno in inovacijsko dejavnost Republike Slovenije (Grant Nos. N1-0154 and P1-0294).
Code availability The code and data for numerical examples are available at https://github.com/borplestenjak/ZGV_Points.
Declarations
Conflict of interests The author has no conflict of interest to declare that are relevant to the content of this article.
Appendix A Generic situation
The following theorem relies on results from algebraic geometry that can be found in, e.g., [12] and [30].
Theorem A.1.
There exist generic sets $\Omega_{3}\subset\Omega_{2}\subset\Omega_{1}\subset(\mathbb{C}^{n\times n})^{3}$ such that:
  1. 1)
    For all $(A,B,C)\in\Omega_{1}$, the matrices $B$ and $C$ are nonsingular, and the GEP $(A+\lambda_{0}B)x+\mu Cx=0$ has a multiple eigenvalue in $\mu$ for only finitely many $\lambda_{0}\in\mathbb{C}$.
  2. 2)
    For all $(A,B,C)\in\Omega_{2}$, the bivariate pencil $A+\lambda B+\mu C$ has exactly n(n1)n(n-1) distinct 2D points.
  3. 3)
    For all $(A,B,C)\in\Omega_{3}$, if $(\lambda_{0},\mu_{0})\in\mathbb{C}^{2}$ is a 2D point of $A+\lambda B+\mu C$, then $(\lambda_{0},\mu_{0})$ is a ZGV point and $\lambda_{0}$ is a double eigenvalue in $\lambda$ of the GEP $(A+\mu_{0}C)x+\lambda Bx=0$.
Proof.
Let $f(\lambda,\mu,A,B,C):=\det(A+\lambda B+\mu C)$.
  1. 1)
    We consider $f$ as a polynomial in $\mu$ whose coefficients are polynomials in $\lambda$ and the entries of $A,B,C$. Then $f$ has a multiple root if the resultant $\mathrm{Res}_{\mu}(f,f_{\mu})$ of $f$ and its derivative $f_{\mu}=\frac{\partial f}{\partial\mu}$ is zero. Since $\mathrm{Res}_{\mu}(f,f_{\mu})$ is a polynomial in $\lambda$ whose coefficients are polynomials in the entries of $A,B,C$, this happens for only finitely many $\lambda$ unless $\mathrm{Res}_{\mu}(f,f_{\mu})\equiv 0$. We can take $\Omega_{1}=(\mathbb{C}^{n\times n})^{3}\backslash(S_{1,a}\cup S_{1,b}\cup S_{1,c})$, where $S_{1,a}=\{(A,B,C):\ \det(B)=0\}$, $S_{1,b}=\{(A,B,C):\ \det(C)=0\}$, and $S_{1,c}=\{(A,B,C):\ \mathrm{Res}_{\mu}(f,f_{\mu})\equiv 0\}$ are algebraic sets and thus $\Omega_{1}$ is generic.
  2. 2)
    Let $(A,B,C)\in\Omega_{1}$. Then $B$ and $C$ are nonsingular and, by Lemma 2.4, $(\lambda_{0},\mu_{0})\in\mathbb{C}^{2}$ is a 2D point if and only if $\lambda_{0}$ is a multiple eigenvalue of the GEP $(A+\mu_{0}C)x+\lambda Bx=0$. If we now consider $f$ as a polynomial in $\lambda$, then $g(\mu,A,B,C):=\mathrm{Res}_{\lambda}(f,f_{\lambda})$, where $f_{\lambda}=\frac{\partial f}{\partial\lambda}$, is a polynomial in $\mu$ of degree n(n1)n(n-1) unless the leading coefficient, which is a polynomial in the entries of $A,B,C$, vanishes. We introduce algebraic sets $S_{2,a}=\{(A,B,C):\ \textrm{coefficient of}\ \mu^{n(n-1)}\ \textrm{in}\ g\ \textrm{is zero}\}$ and $S_{2,b}=\{(A,B,C):\ \mathrm{Res}_{\mu}(g,g_{\mu})=0\}$, and take $\Omega_{2}=\Omega_{1}\backslash(S_{2,a}\cup S_{2,b})$.
    Then $\Omega_{2}$ is generic and for each $(A,B,C)\in\Omega_{2}$ the polynomial $g$ has n(n1)n(n-1) distinct roots. To each root $\mu_{0}$ there corresponds at least one 2D point $(\lambda_{0},\mu_{0})$, where $\lambda_{0}$ is a multiple eigenvalue of the GEP $(A+\mu_{0}C)x+\lambda Bx=0$. Since 2D points are the roots of the system of polynomials $f(\lambda,\mu,A,B,C)=0$ and $f_{\lambda}(\lambda,\mu,A,B,C)=0$, which cannot have more than n(n1)n(n-1) roots by Bézout’s theorem, it follows that $A+\lambda B+\mu C$ has exactly n(n1)n(n-1) distinct 2D points.
  3. 3)
    We observe that $f$ as a polynomial in $\lambda$ has a root of multiplicity three or higher if in addition to $\mathrm{Res}_{\lambda}(f,f_{\lambda})$ also its first subresultant, which is again a polynomial in $\mu$ and the entries of $A,B,C$, is zero, i.e., $\mathrm{Res}_{\lambda}(f,f_{\lambda})=\mathrm{Sres}_{1,\lambda}(f,f_{\lambda})=0$. This happens only if $\mathrm{Res}_{\mu}\big(\mathrm{Res}_{\lambda}(f,f_{\lambda}),\mathrm{Sres}_{1,\lambda}(f,f_{\lambda})\big)=0$, which is a polynomial condition in the entries of $A,B,C$ that we use to define the algebraic set $S_{3,a}$.
    To exclude the possibility of 2D points $(\lambda_{0},\mu_{0})$ such that $\mu_{0}$ is a multiple eigenvalue of the GEP $(A+\lambda_{0}B)x+\mu Cx=0$, i.e., $(\lambda_{0},\mu_{0})$ is a common root of $f,f_{\lambda}$, and $f_{\mu}$, we introduce the algebraic set $S_{3,b}$ in $\mathbb{C}\times\mathbb{C}\times(\mathbb{C}^{n\times n})^{3}$ of all $(\lambda,\mu,A,B,C)$ such that $f(\lambda,\mu,A,B,C)=0$, $f_{\lambda}(\lambda,\mu,A,B,C)=0$, and $f_{\mu}(\lambda,\mu,A,B,C)=0$. The Zariski closure $\overline{\Pi(S_{3,b})}$ of its projection $\Pi(S_{3,b})$ onto $(\mathbb{C}^{n\times n})^{3}$ is algebraic. To show that $\overline{\Pi(S_{3,b})}$ is proper, we need $(A_{0},B_{0},C_{0})\in\Omega_{2}$ such that $f_{\mu}(\lambda,\mu,A_{0},B_{0},C_{0})\neq 0$ for all 2D points of $A_{0}+\lambda B_{0}+\mu C_{0}$.
    Due to Dixon’s result on determinantal representations [7], there exist $n\times n$ matrices $A_{0},B_{0},C_{0}$ such that
    $f(\lambda,\mu,A_{0},B_{0},C_{0})=\lambda^{n}-\lambda-\mu^{n}+1.$ (50)
    We see that $f_{\lambda}(\lambda,\mu)=n\lambda^{n-1}-1$ vanishes when $\lambda^{n-1}=\frac{1}{n}$, which gives n1n-1 simple solutions $\lambda$. By inserting such $\lambda$ in (50), we see that corresponding $\mu$ satisfy the equation $\mu^{n}=1-\frac{n-1}{n}\lambda$, which has $n$ nonzero simple solutions. This way we obtain n(n1)n(n-1) distinct 2D points. Since $f_{\mu}(\lambda,\mu)=-n\mu^{n-1}$ and all 2D points have nonzero $\mu$, it follows that $f_{\mu}$ is nonzero for all 2D points of $A_{0}+\lambda B_{0}+\mu C_{0}$.
    Then, by continuity and implicit function theorem, $f_{\mu}$ is nonzero for all 2D points of the bivariate pencil $A+\lambda B+\mu C$ for all $(A,B,C)$ in a sufficiently small open Euclidean neighborhood of $(A_{0},B_{0},C_{0})$. This shows that $(A_{0},B_{0},C_{0})$ does not lie in the Zariski closure of $\Pi(S_{3,b})$, therefore $\overline{\Pi(S_{3,b})}$ is proper and we can take $\Omega_{3}=\Omega_{2}\backslash(S_{3,a}\cup\overline{\Pi(S_{3,b})})$.
Note that matrices $(A,B,C)\in\Omega_{1}$ satisfy the assumptions of Theorem 4.1. If $(A,B,C)\in\Omega_{3}$, then all 2D points are ZGV points, i.e., of type a) from Remark 2.5, which is the generic case.